### v1

# Antonio Raiola

Abstract
While the literature on grouped patterns in panel data analysis has received significant attention, little to no results are available on testing for their presence. We propose using existing tools for testing slope homogeneity in panels for this purpose. We highlight the key advantages and limitations of the available testing frameworks under a sequence of doubly local alternatives, where slopes are divided into dominant and remainder groups, with the size of the remainder groups and the slopes differences shrinking at a certain rate as the sample size increases. A Monte Carlo study corroborate our theoretical findings.
JEL classification: C13, C23, C33
Keywords: Panel Data; Grouped Patterns; Slope Heterogeneity.
1 Introduction
Driven by empirical demand, the econometric literature on grouped structure in panel data has received significant attention in both econometrics and economics. Most contributions to date focus on identification, estimation, and inference for individual group membership and group-specific parameters; see, e.g., Bonhomme and Manresa (2015), Bester and Hansen (2016), Su et al. (2016), Lu and Su (2017), and Dzemski and Okui (2024). Before estimating such structures, an applied researcher typically faces a more basic diagnostic question: is there evidence of any grouped structure in the first place? This paper propose a simple and readily implementable solution. We repurpose well-known test for slope homogeneity as diagnostics for the presence of a grouped structure, without estimating group membership or the number of groups. We show that these tests provide exactly the needed diagnostic when evaluated under appropriately designed sequences of local alternatives that encode grouped patterns, and we develop theory that clarifies when and why these tests have nontrivial power.
Our approach is guided by practicality: we rely on well-known tools that are readily available in standard software. The literature on testing slope homogeneity is long-standing. A seminal starting point is Swamy (1970)’s a test statistic, based on deviations of individual slope estimates from the pooled estimator; refinements such as Pesaran and Yamagata (2008) remain among the most widely used tools for panel data models. In parallel, several LM tests have been developed for random-coefficient models, where the null of slope homogeneity is equivalent to the variance of the random coefficients being zero. Building on Breusch and Pagan (1979) (see also Hsiao 2003, Sec. 6.2.2.4), Juhl and Lugovskyy (2014) propose LM statistics for zero variance, and Breitung et al. (2016) further extend the methodology to empirically relevant settings. Other LM tests assess whether the regressors have predictive power for the restricted residuals under the null of slope homogeneity–Su and Chen (2013) in models with interactive fixed effects and Lu and Su (2017), originally developed to determine the number of groups.
We analyze the power of established test statistics—Pesaran and Yamagata (2008), Su and Chen (2013), and the LM-type test of Breitung et al. (2016)—under a sequence of doubly local alternatives designed to capture grouped patterns. These alternatives are doubly local in the sense that (i) the population is split into a dominant group and a remainder set of units (possibly partitioned into several groups), and (ii) the slope differences across groups shrink at a controlled rate. This design allows us to characterize how power depends jointly on the convergence rate of group-specific differences and the relative size of the remainder group(s). We consider this setup across several scenarios: two groups; multiple groups, where the number of groups is either fixed or increases with the cross-sectional dimension $N$; and both large-$N$, large-$T$ panels and large-$N$, fixed-$T$ panels, where $T$ denotes the time dimension.
Our analysis yields several findings. We derive the boundaries that characterize when grouped patterns are detectable by linking the rate at which slope differences vanish, $\gamma$, to the size of the remainder group(s), $M$. In large-$N$, large-$T$ panels, the boundary is governed by $\gamma\asymp M^{1/2}T^{1/2}/N^{1/4}$. This result nests the classical $N^{1/4}\sqrt{T}$ rate—when $M/N\to c>0$—reported in Pesaran and Yamagata (2008), Su and Chen (2013), and Breitung et al. (2016), derived under analogous Pitman-drift sequences. Importantly, the tests remain informative even if $M/N\to 0$, provided $MT\to\infty$ and the boundary condition above is satisfied, making sparse grouped alternatives detectable as well. When $T$ is fixed, detectability requires $M$ to be on the order of $\sqrt{N}$ or larger; below this threshold, grouped heterogeneity is undetectable. Finally, if the number of groups increases with $N$, detectability depends on the total number of units in the remainder groups, rather than on the sheer count of groups. A detailed analysis of each scenario is presented in Section 4. Monte Carlo evidence in Section 5 confirm our theoretical findings.
The rest of the paper is structured as follows. Section 2 introduces the model and notation. Section 3 presents the test statistics. Section 4 develops the local-power analysis under doubly local alternatives for the large-$T$ and fixed-$T$ cases, including multiple-group settings. Proofs are relegated to the appendix.
2 Preliminaries
Consider the linear panel data model
$y_{i}=x_{i}^{\prime}\beta_{g(i)}+\epsilon_{i},\quad i=1,\ldots,N,$ (1)
where $y_{i}=(y_{i1},\ldots,y_{iT})^{\prime}$ is the $T$-dimensional vector of dependent variables for unit $i$; $x_{i}=(x_{i1},\ldots,x_{iT})^{\prime}$ is the $T\times K$ design matrix of regressors for unit $i$; and $\epsilon_{i}=(\epsilon_{i1},\ldots,\epsilon_{iT})^{\prime}$ is the corresponding vector of idiosyncratic errors. The slope parameter, $\beta_{g(i)}$, is group-specific, with unknown group membership determined by the mapping $g:\mathcal{N}\to\{G_{1},\ldots,G_{P}\}$, where $\mathcal{N}=\{1,\ldots,N\}$. The function $g$ assigns unit $i$ to one of the $P$ groups partitioning the sample; that is, $G_{p}\cap G_{q}=\emptyset$ for all $p\neq q$, and $\bigcup_{p=1}^{P}G_{p}=\mathcal{N}$.
We are interested in testing for the presence of a grouped structure in the slopes by means of existing slope homogeneity tests. Specifically, we study the local power of such tests against alternatives in which the slope coefficients differ across groups of individuals. The null hypothesis of identical slopes is formulated as,
$H_{0}:\beta_{g(i)}=\beta,\text{ for all }i\in\mathcal{N},$
versus the alternative that there exists a group of individuals with different slopes,
$\G.H_{1}:\exists\ G\subset\{1,...,N\}:\;G\neq\emptyset\;\text{and}\;\beta_{g(i)}\neq\beta_{g(j)}\;\text{for}\;i\in G\;\text{and}\;j\in\mathcal{N}\backslash$ G.
As discussed in the introduction, a variety of well-established tests for slope homogeneity are available in the literature. To analyze their asymptotic behavior within a unified framework, we adopt a common set of assumptions sufficiently broad to capture their essential features.
Throughout, $C$ denotes a generic constant that may vary across expressions, and #G\#G denotes the cardinality of a generic set $G\subseteq\mathcal{N}$.
Assumption 1 (Errors)
$(a)$ $\epsilon_{it}|x_{i}\sim IID(0,\sigma^{2}_{i})$, where $\max_{i}\sigma^{2}_{i}<C<\infty$, $\min_{i}\sigma^{2}_{i}>0$, with $\epsilon_{it}$ and $\epsilon_{js}$ indipendently distributed for $i\neq j$ and $t\neq s$; $(b)$ $\mathbb{E}\left[\epsilon_{it}^{6}|x_{i}\right]<C<\infty$.
Assumption 2 (Regressors)
For all $i\in\mathcal{N}$: $(a)$ $\mathbb{E}|x_{it,k}|^{6}<C<\infty$, for all t=1,,Tt=1,...,T and k=1,,Kk=1,...,K; $(b)$ the probability limit $S_{i}=\mathop{\rm plim}_{T\to\infty}T^{-1}x_{i}^{\prime}x_{i}$, is a positive definite, non-stochastic and finite matrix; (c) for any subset $G\subseteq\mathcal{N}$ such that $G\to\infty$, the probability limit $S^{(G)}=\mathop{\rm plim}_{N,T\to\infty}(\#G)^{-1}\sum_{i\in G}S_{i}$ is a positive definite, non-stochastic and finite matrix.
In particular, taking $G=\mathcal{N}$, $S=S^{(\mathcal{N})}$ is a positive definite non-stochastic matrix.
Assumption 1(a), corresponding to Assumption 1(i)–(ii) in Pesaran and Yamagata (2008), rules out serial correlation, and cross-sectional dependence. Together with model (1), which excludes endogenous components in the error term, they provide a simplified framework for presenting our main results in a unified manner. We emphasize, however, that the tests considered here have been developed or extended to accommodate more general error structures (see original papers as well as Blomquist and Westerlund 2013; Ando and Bai 2015). In all such cases, the main conclusions of our analysis remain valid.
We focus here on three representative and widely used slope homogeneity tests: (i) the dispersion statistic $\Delta$ from Pesaran and Yamagata (2008) (hereafter, PY); (ii) the residual-based LM test $\mathcal{J}$ from Su and Chen (2013) (SC); and (iii) the LM-type variance test $LM$ from Breitung et al. (2016) (BRS). The next section provides the necessary details for each of the test statistics.
3 Test Statistics
The $\Delta$ test of PY refines the Swamy (1970) test, originally designed for testing slope homogeneity in large $T$, fixed $N$, scenarios. Both tests compare individual slope estimates, $\hat{\beta}_{i}=(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}y_{i}$, with the weighted least squares (WLS) estimator $\hat{\beta}_{WLS}(\bar{\sigma}_{i}^{2})=\bigl(\sum_{i=1}^{N}x_{i}^{\prime}x_{i}/\bar{\sigma}_{i}^{2}\bigr)^{-1}\bigl(\sum_{i=1}^{N}x_{i}^{\prime}y_{i}/\bar{\sigma}_{i}^{2}\bigr)$, using some estimator $\bar{\sigma}^{2}_{i}$ of $\sigma^{2}_{i}$. The original Swamy statistic is $\mathcal{S}_{SW}=\mathcal{S}(\hat{\sigma}^{2}_{i})$, where
$\mathcal{S}(\bar{\sigma}^{2}_{i})=\sum_{i=1}^{N}\left(\hat{\beta}_{i}-\hat{\beta}_{WLS}(\hat{\sigma}^{2}_{i})\right)^{\prime}\left(\frac{x_{i}^{\prime}x_{i}}{\bar{\sigma}^{2}_{i}}\right)\left(\hat{\beta}_{i}-\hat{\beta}_{WLS}(\hat{\sigma}^{2}_{i})\right),$
and $\hat{\sigma}^{2}_{i}=(T-K-1)^{-1}(y_{i}-x_{i}\hat{\beta}_{i})^{\prime}(y_{i}-x_{i}\hat{\beta}_{i})$. This statistic has a $\chi^{2}$ null distribution with K(N1)K(N-1) degrees of freedom when $T\to\infty$ and $N$ is fixed. For the large-$N$, large-$T$ framework, PY propose to replace $\hat{\sigma}^{2}_{i}$ with
$\widetilde{\sigma}^{2}_{i}=\frac{\hat{\epsilon}_{i}^{\prime}\hat{\epsilon}_{i}}{T-K-1},$
where $\hat{\epsilon}_{i}=(\hat{\epsilon}_{i1},...,\hat{\epsilon}_{iT})^{\prime}$, $\hat{\epsilon}_{it}=y_{it}-x_{it}^{\prime}\hat{\beta}_{LS}$, and $\hat{\beta}_{LS}=(\sum_{i=1}^{N}x_{i}^{\prime}x_{i})^{-1}\sum_{i=1}^{N}x_{i}^{\prime}y_{i}$ - the pooled least square estimator. The resulting dispersion statistic, $\mathcal{S}_{PY}=\mathcal{S}(\tilde{\sigma}^{2}_{i})$, after centering and standardization, converges to the standard normal distribution under $H_{0}$,
$\Delta=\frac{\mathcal{S}_{PY}-NK}{\sqrt{2NK}}\overset{d}{\rightarrow}\mathcal{N}(0,1),\quad\text{ as }\quad N,T\to\infty.$ (2)
The residual-based $\mathcal{J}$ test. An alternative approach is the LM test of SC, [^2] With additive fixed effects (or none, as in the present case), $\mathcal{J}$ is numerically equivalent to the test of Lu and Su (2017) when applied to a single-group case under the null. based on the idea that, under $H_{0}$ and with predetermined regressors, $x_{it}$ should not have explanatory power for the residuals in (1). This motivates a test of $H_{0}^{*}:\phi_{i}=0\;\text{for all }i=1,\ldots,N$ in the auxiliary regression
$\hat{\epsilon}_{it}=\phi_{i}^{\prime}x_{it}+\eta_{it},\qquad\text{with}\quad i=1,\ldots\,N,t=1,\ldots,T,$
where $\eta_{it}$ is a mean-zero error term.
Let $h_{i}=x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}$ be the projection matrix with diagonal elements $h_{i,tt}$, and define $\hat{b}_{it}=\hat{S}_{i}^{-1/2}x_{it}$, where $\hat{S}_{i}=T^{-1}x_{i}^{\prime}x_{i}$. The SC test statistic is given by the following centered and standardized $LM$ test statistic, which has a limiting normal distribution under $H_{0}^{*}$:
$\mathcal{J}=\frac{N^{-1/2}LM_{SC}-\hat{B}_{NT}}{\sqrt{\hat{V}_{NT}}}\overset{d}{\rightarrow}\mathcal{N}(0,1),\quad N,T\to\infty,$ (3)
where
$LM_{SC}=\sum_{i=1}^{N}\hat{\epsilon}_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\hat{\epsilon}_{i},$
and $\hat{B}_{NT}=N^{-1/2}\sum_{i=1}^{N}\sum_{t=1}^{T}\hat{\epsilon}_{it}^{2}h_{i,tt}$ and $\hat{V}_{NT}=4T^{-2}N^{-1}\sum_{i=1}^{N}\sum_{t=2}^{T}\left[\hat{\epsilon}_{it}\hat{b}_{it}^{\prime}\sum_{s=1}^{t-1}\hat{b}_{is}\hat{\epsilon}_{is}\right]^{2}$ estimate the mean and variance of $LM_{SC}$ under the null, respectively.
The LM variance test. A third approach reformulates the null hypothesis within the random coefficients model. Specifically, assuming $\beta_{i}=\beta+v_{i}$ with $\mathbb{E}[v_{i}v_{i}^{\prime}|x_{i}]=\text{diag}(\sigma^{2}_{v,1},\dots,\sigma^{2}_{v,K})$, slope homogeneity is equivalent to $H_{0}^{**}:\sigma^{2}_{v,k}=0$, for all $k=1,\dots,K$, which reduces the N(K1)N(K-1) linear restrictions tested by $\Delta$ and $\mathcal{J}$ to $K$.
Building on this idea, BRS develop an LM test valid for panels with large $N$ and either fixed $T$ or large $T$. The statistic is based on the score vector $\hat{s}=(\hat{s}_{1},\ldots,\hat{s}_{K})^{\prime}$, where
$\hat{s}_{k}=\sum_{i=1}^{N}\sum_{t=2}^{T}\sum_{s=1}^{t-1}\hat{\epsilon}_{it}\hat{\epsilon}_{is}x_{it,k}x_{is,k},\;\;k=1,...,K.$
The test statistic is then given by
$LM=\hat{s}^{\prime}\hat{V}^{-1}\hat{s}$, (4)
where $\hat{V}$ has $(k,l)$-th element
$\hat{V}_{kl}=\sum_{i=1}^{N}\sum_{t=2}^{T}\sum_{s=1}^{t-1}\sum_{p=1}^{t-1}\hat{\epsilon}_{it}^{2}x_{it,k}x_{it,l}\,\hat{\epsilon}_{is}\hat{\epsilon}_{ip}\,x_{is,k}x_{ip,l}\quad\text{and}\quad l,k=1,...,K.$
Under $H_{0}^{**}$ and $N\to\infty$, with $T$ either fixed or $T\to\infty$, the test has a $\chi^{2}_{K}$ null distribution.
The three tests, $\Delta$, $\mathcal{J}$, and $LM$, are consistent for $H_{0}$ and have well-established (local) power properties. However, none has been explicitly studied in the context of local power against grouped slope structures, which is the focus of the next section.
4 Asymptotic Local Power Analysis
We study the local power of $\Delta$, $\mathcal{J}$, and $LM$ tests under grouped alternatives, focusing on (i) how large slope differences must be to be detected and (ii) how large the group must be for detection. To this end, we introduce a sequence of doubly local alternatives that extends the classical Pitman framework and integrates these two aspects. We first present the results for the case where the sample is partitioned into two groups, $\{G_{1},G_{2}\}$, with $N,T\to\infty$. Group $G_{1}$ is treated as the “dominant” group and $G_{2}$ as the “alternative” one, containing the remaining individuals. We then extend these results to panels with $N\to\infty$ and fixed $T$, as well as to settings with more than two groups.
Formally, we define
$H_{1,n}:\beta_{g(i)}=\begin{cases}\beta,&\text{if }i\in G_{1},\\ \beta+\lambda_{2}/\gamma,&\text{if }i\in G_{2},\end{cases}$
where $\lambda_{2}\neq 0$ is a $K\times 1$ vector, and $\{G_{1},G_{2}\}$ form a partition of $\mathcal{N}$. The number of individuals in each group is
$G_{2}=M,\qquad\#G_{1}=N-M,$
where $M\to\infty$ as $N\to\infty$ and $\gamma=\gamma(N,T,M)$ diverges with $N$, $T$, and $M$.
Allowing the scaling factor $\gamma$ to depend on $M$ as well as on $N$ and $T$ reflects that our proposed local alternatives control for both the slope difference and the size of the heterogeneous group. This extension is essential: the relative magnitude of $M$ with respect to $N$ and $T$ is crucial in determining when and why slope homogeneity tests can detect group-based heterogeneity. Theorem 1 formalizes this statement.
Theorem 1
Under Assumptions 1, 2, and $H_{1,n}$, as $M,N,T\to\infty$ and $M/N\to{m_{0}}\in[0,1)$, the following holds:
(PY test)
Additionally assume $\mathbb{E}\left[\epsilon_{it}^{9}|x_{i}\right]<C<\infty$ and $\sqrt{N}/T^{2}\to 0$, then,
$\Delta\overset{d}{\rightarrow}\begin{cases}\mathcal{N}(0,1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to 0,\\ \mathcal{N}(\delta_{\Delta},1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to\infty.\end{cases},$
where
${\delta}_{\Delta}=\frac{c^{2}}{\sqrt{2K}}\lambda_{2}^{\prime}\left(Q^{(G_{2})}-m_{0}Q^{(G_{2})}Q^{-1}Q^{(G_{2})}\right)\lambda_{2},$
with $Q^{(G)}=\mathop{\rm plim}\limits_{\#G,T\to\infty}(\#GT)^{-1}\sum_{i\in G}x_{i}^{\prime}x_{i}/\sigma^{2}_{i}$, and $Q=Q^{(\mathcal{N})}$.
(SC test)
$\mathcal{J}\overset{d}{\rightarrow}\begin{cases}\mathcal{N}(0,1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to 0,\\ \mathcal{N}(\delta_{\mathcal{J}},1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to\infty.\end{cases},$
where
${\delta}_{J}=\frac{c^{2}}{\sqrt{V_{0}}}\lambda_{2}^{\prime}\left(W^{(G_{2})}+m_{0}(S^{(G_{2})}S^{-1}WS^{-1}S^{(G_{2})}-2W^{(G_{2})}S^{-1}S^{(G_{2})})\right)\lambda_{2},$
with $W^{(G)}=\mathop{\rm plim}\limits_{\#G,T\to\infty}(\#GT)^{-1}\sum_{i\in G}\sum_{t=1}^{T}{x_{it}x_{it}^{\prime}}(1-h_{i,tt})$, and $W=W^{(\mathcal{N})}$. Here $V_{0}=\mathop{\rm plim}_{N,T\to\infty}4T^{-2}N^{-1}\sum_{i=1}^{N}\sum_{t=2}^{T}\mathbb{E}\left[\left(\epsilon_{it}b_{it}^{\prime}\sum_{s=1}^{t-1}b_{is}\epsilon_{is}\right)^{2}\right]$, with $b_{it}=S_{i}x_{it}$.
(BRS test)
$LM\overset{d}{\rightarrow}\begin{cases}\chi_{K}^{2},&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to 0,\\ \chi_{K}^{2}\left(\delta_{LM}^{\prime}\Psi^{-1}\delta_{LM}\right),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to\infty.\end{cases},$
where $\delta_{LM}$ is a $K\times 1$ vector, with $k$-th element
$\delta_{LM,k}=c^{2}\left(\lambda_{2}^{\prime}\Omega_{k}^{(G_{2})}\lambda_{2}+m_{0}\lambda_{2}^{\prime}\left(S^{(G_{2})}S^{-1}\Omega_{k}S^{-1}S^{(G_{2})}-2S^{(G_{2})}S^{-1}\Omega_{k}^{(G_{2})}\right)\lambda_{2}\right)$
with $\Omega_{k}^{(G)}=\mathop{\rm plim}_{\#G,T\rightarrow\infty}\;\frac{1}{\#GT^{2}}\sum_{i\in G}\sum_{s<t}x_{it,k}x_{it}x_{is,k}x_{is}^{\prime}$, and $\Omega_{k}=\Omega_{k}^{(\mathcal{N})}$. Here, $\Psi$} is a $K\times K$ matrix with ${\Psi}_{k,l}=\underset{N,T\rightarrow\infty}{\text{plim}}\;(NT^{2})^{-1}\sum_{i=1}^{N}\sum_{s<t}\sigma^{4}_{i}x_{it,k}x_{it,l}x_{is,k}x_{is,l}$.
Theorem 1 highlights a fundamental trade-off between the rate at which slope differences vanish and the size of the alternative group. Its key implication is the rate
$\gamma\asymp\frac{M^{1/2}T^{1/2}}{N^{1/4}},\qquad M,N,T\to\infty,$ (5)
which controls the transition between trivial and nontrivial power. Specifically, if $\gamma$ diverges faster than the rate in (5), all three tests have trivial power; otherwise they have power to detect the grouped alternative.
It is worth emphasizing when the result in (5) reproduces earlier findings and when it yields genuinely new insights. If the fraction of alternative units remains bounded away from zero, $m_{0}>0$ (so that $M\asymp N$), the critical rate in (5) reduces to the classical order $O(\sqrt{T}N^{1/4})$. In this case, our framework nests the local power behavior established by PY, SC, and BRS, providing a consistency check with existing results. The theorem further derives noncentrality parameters that differ from those in the original papers and clarify the distinct role of group-based heterogeneity in local power.
Further novel implications arise when the alternative group is small relative to the dominant one, i.e., when $m_{0}=0$. In this case, the rate in (5) is slower than $O(\sqrt{T}N^{1/4})$; nevertheless, the tests remain informative as long as $M$ diverges with $N$, achieving nontrivial power even when only a small fraction of units (i.e., $M/N\to 0$) display distinct slopes. This contrasts with the conventional intuition in the grouped-patterns literature, where heterogeneous slopes are typically assumed to affect a non-negligible share of the sample (e.g., Bonhomme and Manresa, 2015; Su et al., 2016; Lu and Su, 2017). These results highlight the usefulness of slope homogeneity tests in empirical contexts where the presence of small, systematically different subgroups matters for inference and policy.
4.1 Panels with $N\to\infty$ and $T$-fixed
When $T$ is fixed and $N$ grows, the only available slope homogeneity test among those considered is the LM-type test developed by BRS. The restrictions on the regressors can be relaxed, as in Assumption 3 in the Appendix, which also contains the corresponding version of Theorem 1. In this scenario, the boundary between trivial and nontrivial power is governed by the rate
$\gamma\asymp\frac{\sqrt{M}}{N^{1/4}},\qquad M,N\to\infty,$ (6)
the analogue of (5) in this setting.
When the size of the alternative group is proportional to the total cross-section, $m_{0}>0$ (or $M\!\asymp\!N$), the rate in (6) reduces to $\gamma\!\asymp\!N^{1/4}$, which coincides with the local-power behavior established in BRS. More interestingly, the rate in (6) marks the boundary at which the test begins to exhibit power. As long as $M$ diverges at least at the rate $N^{1/2}$, the $LM$ test retains nontrivial local power even when $M/N\!\to\!0$. Essentially, this result identifies the minimal detectable group size as being of order $N^{1/2}$: below this threshold, grouped heterogeneity is undetectable.
4.2 Panels with increasing number of groups
The main message of our analysis will remain valid even if we allow the number of groups to grow with $N$. Consider
$H_{2,n}:\beta_{g(i)}=\begin{cases}\beta,&\mbox{if }i\in G_{1},\\ \beta+\lambda_{p}/\gamma,&\mbox{if }i\in G_{p}\mbox{ and }p=2,...,P,\end{cases}$
where $\{G_{1},...,G_{P}\}$ forms a partition of $\mathcal{N}$, with $\lambda_{p}\neq\lambda_{q}$ for any $p,q\in\{1,...,P\}$ such that $p\neq q$. We let $M_{p}=\#G_{p}$ and $\mathbb{M}=\sum_{p=2}^{P}M_{p}$.
The analysis of local power when $P\to\infty$ as $N\to\infty$ is complicated by the fact that the number of individuals in each group $\{G_{2},...,G_{P}\}$ can also be a function of $N$. Hence, the combination of both number of groups and number of individuals in each group plays a key role to obtain the expansion rates under which the test statistics have nontrivial power.
Corollary 1
Let Assumptions 1, 2, and the sequence of local alternatives $H_{2,n}$ hold. If $\mathbb{M}\to\infty$ as $N\to\infty$, then the result of Theorem 1 holds true by replacing $M$ with $\mathbb{M}$, and the non-centrality parameters, $\delta_{\Delta}$, $\delta_{\mathcal{J}}$, and $\delta_{LM,k}$ with,
$\lim_{N,T\to\infty}\frac{1}{\sqrt{2K}}\left(\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}\left(Q^{(G_{p})}\lambda_{p}\right)-\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}\left(Q^{(G_{p})}Q^{-1}Q^{(G_{q})}\right)\lambda_{q}\right),$
$\lim_{N,T\to\infty}\frac{1}{\sqrt{V_{0}}}\left(\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}W^{(G_{p})}\lambda_{p}+\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}\left[S^{(G_{p})}S^{-1}WS^{-1}S^{(G_{q})}-2W^{(G_{p})}S^{-1}S^{(G_{q})}\right]\lambda_{q}\right)$
and,
$\lim_{N,T\to\infty}\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}U_{k}^{(G_{p})}\lambda_{p}+\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}^{\prime}\left(S^{(G_{p})}S^{-1}\Omega_{k}S^{-1}S^{(G_{q})}-2S^{(G_{p})}S^{-1}\Omega_{k}^{(G_{q})}\right)\lambda_{q},$
respectively, where for all p=2,,Pp=2,...,P, $M_{p}/N\to m_{p}\in[0,1)$ and $M_{p}T/(\gamma\sqrt{N})\to c_{p}^{2}\in[0,c^{2}]$, with $c=\mathbb{M}T/(\sqrt{N}\gamma^{2})$.
Corollary 1 generalize the group structure in Theorem 1. To see that Corollary 1 nests previous results assume that $P$ is fixed and $M^{\ast}=\max_{2\leq p\leq P}M_{p}\to\infty$ as $N\to\infty$. Then by replacing $M=M^{\ast}$ the results in Theorem 1 follows. Further, this corollary also nests the case when $P\to\infty$ as $N,T\to\infty$ and the number of individuals in groups $\{G_{2},...,G_{P}\}$ is fixed and bounded (i.e., $\sup_{2\leq p\leq P}M_{p}=\kappa<\infty$). Then, the expansion rates is given by $PT/(\sqrt{N}\gamma^{2})\to c>0$ with $N,T\to\infty$. Finally, in the general case when both $P\to\infty$ and $M_{p}\to\infty$ for at least one p=2,,Pp=2,...,P, the corresponding expansion rate will be driven by the overall “volume” of individual in $\{G_{2},...,G_{P}\}$, defined by $\mathbb{M}$.
5 Monte Carlo Study
We evaluate the finite-sample performance of the tests in terms of their power, relating to and verifying the theoretical results provided in the main text. We consider the three test statistics ($\Delta$, $LM$, and $\mathcal{J}$) under a data-generating process adapted from PY to fit the context of grouped patterns; i.e., for each $i=1,\ldots,N$ and $t=1,\ldots,T$:
$y_{it}$ $=x_{it}\beta_{i}+\alpha_{i}+\epsilon_{it},$
$x_{it}$ $=\rho_{i}x_{it-1}+\sqrt{1-\rho_{i}^{2}}\,v_{it},$
where $\epsilon_{it}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{i}^{2})$, $v_{it}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{ix}^{2})$, $\alpha_{i}\overset{iid}{\sim}\mathcal{N}(1,1)$, and $\rho_{i}\overset{iid}{\sim}\mathcal{U}(0.05,0.95)$, with $\sigma_{i}^{2}\sim\mathcal{X}^{2}(2)/2$ and $\sigma_{ix}^{2}\sim\mathcal{X}^{2}(1)$. The first 50 observations of $x_{it}$ ($t=-49,\ldots,0$) are discarded to mitigate the effect of initial values. Group-specific slopes are generated as:
$\beta_{i}=\begin{cases}1,&\text{if }i\in G_{1},\\ 1+\lambda(1+e_{p}),&\text{if }i\in G_{p},\end{cases}$
where $p=2,\ldots,P$ and $e_{p}\overset{iid}{\sim}\mathcal{U}(-h,h)$. The first group, $G_{1}$, contains $N-\sum_{p=2}^{P}M_{p}$ individuals, while the remaining P1P-1 groups $\{G_{2},\ldots,G_{P}\}$ represent alternative groups of size $M_{p}$ each. The parameters $\lambda$ and $h$ control the heterogeneity between the dominant and remaining groups, and across alternative groups, respectively.
We consider two simulation designs to examine how the cardinality of the alternative group and the magnitude of slope differences affect statistical power. Each design considers $(N,T)=(100,10)$ and $(100,100)$, corresponding to panels of small and large time dimensions, respectively:
  • (1)
    Two-group design: $P=2$, $h=0$, varying the size of the alternative group $M_{2}$ for $\lambda\in\{0.1,0.3,0.5\}$.
  • (2)
    Multiple-group design: $P\in\{3,5,10\}$, where alternative groups size are randomly assigned subject to $\sum_{p=2}^{P}M_{p}=\mathbb{M}$, with varying $\mathbb{M}$. Here, $\lambda=0.1$ and $h\in\{0.2\}$, focusing on the effect of the overall mass of alternative groups $\mathbb{M}$.
The data-generating process includes individual constants, removed through appropriate demeaning procedures for each test statistic. Specifically, for the $LM$ test, we use orthogonal deviations as in Arellano and Bover (1995), following Remark 5.1 in BRS. For the $\Delta$ and $\mathcal{J}$ tests, we apply the classical demeaning procedure, pre-multiplying $y_{i}$ and $x_{i}$ by $M_{0}=I_{T}-T^{-1}\textbf{i}_{T}\textbf{i}_{T}^{\prime}$, where $\textbf{i}_{T}$ is a $T\times 1$ vector of ones, as suggested in PY and Lu and Su (2017). All simulations report rejection rates at the 5% significance level, based on 1,000 Monte Carlo replications. The results are presented in Figures 1 and 2, respectively.
Two Groups $N=100$, $T=10$
Two Groups $N=100$, $T=100$
Figure 1: Power curves of slope homogeneity tests with two groups.
The two-group scenario results (Figure 1) closely mirror our theoretical findings in Theorem 1. The rejection rates increase monotonically with the alternative group size $M_{2}$ and with the slope deviation parameter $\lambda$. All tests tend to have a power close to trivial when $M_{2}$ is smaller than $\sqrt{N}$, but the power increases significantly after this threshold and when $T$ grows.
In the multiple-group experiment (Figure 2), the decisive role is played by the number of individuals in the alternative groups, $\mathbb{M}$. Tests exhibit substantial power whenever $\mathbb{M}$ grows at least proportionally to $\sqrt{N}$, consistent with Corollary 1. Notice that the tests power decreases with larger $P$, but this has only a secondary impact once $\mathbb{M}$ and/or $T$ is large enough.
$P$ Groups $N=100$, $T=10$
$P$ Groups $N=100$, $T=100$
Figure 2: Power curves of slope homogeneity tests with $P$ groups.
6 Conclusions
We propose repurposing existing slope-homogeneity tests to detect grouped patterns in panels. The power of the tests is evaluated under a sequence of doubly local alternatives, linking the degree of heterogeneity across groups and each group size to its impact on test performance. Our analysis shows that slope-homogeneity tests are effective in detecting such structures, including sparse groups whose cardinality is negligible relative to the cross-sectional size. Monte Carlo simulations support similar conclusions in finite samples.
Appendix
Appendix A Proofs
We derive the asymptotic distribution of the three type of tests in Theorems 1 under $H_{1,n}$, in the large panel setting, where $N,T\to\infty$. The rates stated in Corollaries 1 and 2 (at the end of the Appendix) follow immediately by setting $M=K_{N}$ and $T=1$, respectively. Additional remarks on the derivation of the noncentrality parameters in Corollary 1 are provided at the end of this section.
The proof is based on the decomposition of each test into a term leading the asymptotic distribution of the test, a term that drives the power of the test, and remainder terms, denoted as $R_{NT}$ and $C_{NT}$, that are asymptotically negligible under the conditions of the statements.
Under $H_{1,n}$, we can rewrite model (1) as
$y_{it}=x_{it}^{\prime}\beta_{i}+\epsilon_{it}=x_{it}^{\prime}\beta+x_{it}^{\prime}\lambda_{i}+\epsilon_{it}=x_{it}^{\prime}\beta+u_{it},$
where $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$, $\lambda_{i}=0$ if $i\in G_{1}$ and $\lambda_{i}=\lambda/\gamma$ if $i\in G_{2}$.
Lemma 1
Under $H_{1,n}$ and Assumptions 1, 2,
$\hat{\beta}_{LS}-\beta=O_{p}\left(\frac{M}{N\gamma}+\frac{1}{\sqrt{NT}}\right),\,\mbox{as }N,T\to\infty.$
Proof of Lemma 1.
$\hat{\beta}_{LS}-\beta$ $=$ $\left(\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\right)^{-1}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}\epsilon_{it}+\left(\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\right)^{-1}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\lambda_{i}.$
From Assumption 2(b) we have
$\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}$ $\overset{p}{\to}$ $S\text{ as }N,T\to\infty$
Further, $\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}\epsilon_{it}=O_{p}\left((NT)^{-1/2}\right)$ as $N,T\to\infty$. Finally, from Assumption 2(c)
$\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\lambda_{i}=\frac{1}{\gamma}\sum_{i\in G_{2}}^{N}\sum_{t=1}^{T}{\lambda}x_{it}x_{it}^{\prime}=O_{p}\left(\frac{MT}{\gamma}\right),\mbox{as }N,T\to\infty.$
Proof Theorem 1, BRS Test.
We start by considering the $LM$ test of BRS under $H_{1,n}$. After substituting $\hat{\epsilon}_{it}=u_{it}-x_{it}^{\prime}(\hat{\beta}_{LS}-\beta)$, with $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$, into the score vector, we obtain the following decomposition,
$\hat{s}_{k}=\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\epsilon_{it}\epsilon_{is}x_{it,k}x_{is,k}+\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda_{i}^{\prime}{x}_{it}x_{it,k}x_{is,k}{x}_{is}^{\prime}\lambda_{i}+R_{NT,k}+C_{NT,k},$ (A1)
where the first two terms drive the asymptotic normality and the power of the test, respectively, while $R_{NT,k}$ and $C_{NT,k}$ are remainder terms given by,
$R_{NT,k}$ $=$ $\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}(\hat{\beta}_{LS}-\beta)^{\prime}x_{it}x_{it,k}x_{is,k}x_{is}^{\prime}(\hat{\beta}_{LS}-\beta)-2\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}(\hat{\beta}_{LS}-\beta)^{\prime}x_{it}u_{is}x_{it,k}x_{is,k},$
$C_{NT,k}$ $=$ $\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda_{i}^{\prime}{x}_{it}\epsilon_{is}x_{it,k}x_{is,k}+\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\epsilon_{it}\lambda_{i}^{\prime}{x}_{is}x_{it,k}x_{is,k},$
for k=1,,Kk=1,...,K. After substituting $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$ into $R_{NT}$ and using Lemma 1, Lemma A2 in BRS, together with the fact that $\lambda_{i}$ is non zero for only $M$ elements, we obtain,
$\frac{1}{\sqrt{N}T}R_{NT,k}=O_{p}\left(\frac{M^{2}T}{\gamma^{2}{N}^{3/2}}\right).$
Furthermore, if $M/N\to{m_{0}}>0$ then the rate above reduces to $\frac{1}{\sqrt{N}T}R^{(k)}_{NT}=O_{p}\left(\frac{MT}{\gamma^{2}{N}^{1/2}}\right)$ and as $MT/\sqrt{N}\gamma^{2}\to c>0$ we have
$\frac{1}{\sqrt{N}T}R_{NT,k}$ $=c^{2}m_{0}\lambda^{\prime}\left(S^{(G_{2})}S^{-1}\Omega_{k}S^{-1}S^{(G_{2})}-2S^{(G_{2})}S^{-1}\Omega_{k}^{(G_{2})}\right)\lambda+o_{p}(1).$
By the same arguments,
$\frac{1}{\sqrt{N}T}C_{NT,k}=O_{p}\left(\frac{\sqrt{MT}}{\sqrt{N}\gamma}\right),$
and if $MT/(\sqrt{N}\gamma^{2})\to c\geq 0$ then $\frac{\sqrt{MT}}{\sqrt{N}\gamma}\to 0$ resulting in $\frac{1}{\sqrt{N}T}C_{NT,k}=o_{p}(1)$. Further,
$\frac{1}{\sqrt{N}T}\sum\limits_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda_{i}^{\prime}{x}_{it}x_{it,k}x_{is,k}{x}_{is}^{\prime}\lambda_{i}$ $=\frac{1}{\gamma^{2}\sqrt{N}T}\sum_{i\in G_{2}}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda^{\prime}{x}_{it}x_{it,k}x_{is,k}{x}_{is}^{\prime}\lambda$
$\overset{p}{\rightarrow}\lambda^{\prime}\Omega_{k}^{(G_{2})}\lambda\times\lim\limits_{N,T\to\infty}\frac{MT}{\sqrt{N}\gamma^{2}}.$
Let $s_{k}:=\sum_{i=1}^{N}\sum_{t=2}^{T}\sum_{s=1}^{t-1}\epsilon_{it}\epsilon_{is}x_{it,k}x_{is,k}$ and $s=[s_{1},...,s_{K}]^{\prime}$ then by Theorems 5.1 and Corollary 5.2 in BRS,
$\frac{1}{\sqrt{N}T}s\overset{d}{\rightarrow}N(0,\Psi)\text{ as }N,T\to\infty.$
Finally, following Theorem 6.2 in BRS, we can write each $(k,l)$ element of $\hat{V}$ as,
$\frac{1}{NT^{2}}\hat{V}_{k,l}$ $=\frac{1}{NT^{2}}\sum\limits_{i=1}^{N}\sum\limits_{t=2}^{T}\epsilon_{it}^{2}x_{it,k}x_{it,l}\left(\sum\limits_{s=1}^{t-1}\epsilon_{is}x_{is,k}\right)\left(\sum\limits_{p=1}^{t-1}\epsilon_{ip}x_{ip,k}\right)+O_{p}\left(\frac{MT}{\gamma^{2}N}\right)$
$=\Psi_{k,l}+O_{p}\left(\frac{MT}{\gamma^{2}N}\right).$
Putting all these results together into the expression for the LM test (4) yields the first statement of the theorem.
Proof Theorem 1, PY Test.
We now use the same procedure to prove the statement for the PY test. Under $H_{1,n}$, we have that,
$\left(\hat{\beta}_{i}-\hat{\beta}_{WLS}(\tilde{\sigma}_{i})\right)=\left[\lambda_{i}+\left(\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\frac{x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}\right)-\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}u_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right]$
where $u_{i}=(u_{i1},...,u_{iT})^{\prime}=x_{i}\lambda_{i}+\epsilon_{i}$. Thus, rewrite $S_{PY}$ as,
$S_{PY}=\sum_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}+\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}+R_{NT}+C_{NT},$
where,
$R_{NT}$ $=$ $-\left(\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}\right)-\left(\sum\limits_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}\right),$
$C_{NT}$ $=$ $2\sum_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}-2\left(\sum\limits_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}\right).$
Using that $(NT)^{-1}\sum_{i=1}^{N}(x_{i}^{\prime}x_{i}/\tilde{\sigma}^{2}_{i})=O_{p}(1+T^{-1/2})$ and $(NT)^{-1/2}\sum_{i=1}^{N}(x_{i}^{\prime}\epsilon_{i}/\tilde{\sigma}^{2}_{i})=O_{p}(1+T^{-1/2})$ (see equations A.10, A.11, and Lemma 3 in PY), it follows that,
$\frac{1}{\sqrt{N}}R_{NT}$ $=O_{p}\left(\frac{M^{2}T}{\gamma^{2}{N}^{3/2}}\right).$
Furthermore, if $M/N\to{m_{0}}>0$ then the rate above reduces to $\frac{MT}{\gamma^{2}{N}^{1/2}}$, as $M/N<1$, and if $MT/\sqrt{N}\gamma^{2}\to c>0$ we have
$\frac{1}{\sqrt{N}}R_{NT}=-(c)^{2}m_{0}\lambda^{\prime}Q^{(G_{2})}Q^{-1}Q^{(G_{2})}\lambda.$
By the same arguments
$\frac{1}{\sqrt{N}}C_{NT}$ $=O_{p}\left(\frac{\sqrt{MT}}{\gamma\sqrt{N}}+\frac{M\sqrt{T}}{\gamma N}\right),$
and,
$\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}=\frac{MT}{\sqrt{N}\gamma^{2}}\lambda^{\prime}Q^{(G_{2})}\lambda+O_{p}\left(\frac{M\sqrt{T}}{\sqrt{N}\gamma^{2}}\right).$
Finally, by the same arguments of Theorem 1 and 2 in PY (see appendix A.2, A.3 of PY),
$\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}=\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\tilde{z}_{i}+O_{p}\left(\frac{1}{\sqrt{T}}\right)+O_{p}\left(\frac{1}{\sqrt{N}}\right),$
where $\tilde{z}_{i}=(T-1)\epsilon_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\epsilon_{i}/(\epsilon_{i}^{\prime}\epsilon_{i})$ with $\mathbb{E}\left[\tilde{z}_{i}\right]=K+O(T^{-2})$ and $\operatorname{Var}{\tilde{z}_{i}}=2K+O(T^{-1})$ uniformly over $i$. Then putting all these rates together into expression (2) for PY test yields
$\Delta$ $=\sqrt{N}\left(\frac{N^{-1}S_{PY}-K}{\sqrt{2K}}\right)=\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\left(\frac{\tilde{z}_{i}^{*}-K}{\sqrt{2K}}\right)+O\left(\frac{\sqrt{N}}{T^{2}}\right)$
$+\frac{MT}{\sqrt{N}\gamma^{2}}\lambda^{\prime}\left(\frac{1}{MT}\sum_{i\in G_{2}}\frac{x_{i}^{\prime}x_{i}}{\sigma_{i}^{2}}\right)\lambda\frac{1}{\sqrt{2K}}+\frac{MT}{\sqrt{N}\gamma^{2}}O_{p}\left(\frac{M}{N}\right)\frac{1}{\sqrt{2K}}+o_{p}(1),$
where $N^{-1/2}\sum_{i=1}^{N}(\tilde{z}_{i}-K)/\sqrt{2K}\xrightarrow{d}N(0,1)$ (appendix A.4 of PY).
The statement in Theorem 1 follows immediately by considering different scenarios for the rate $\frac{MT}{\sqrt{N}\gamma^{2}}$. ∎
Proof Theorem 1, SC Test.
After substituting $\hat{\epsilon}_{i}=\epsilon_{i}+x_{i}\lambda_{i}-x_{i}(\hat{\beta}_{LS}-\beta)$ into $LM_{SC}-\sqrt{N}\hat{B}_{NT}$ we obtain,
$LM_{SC}-\sqrt{N}\hat{B}_{NT}=\sum_{i=1}^{N}\epsilon_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})\epsilon_{i}+\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}\lambda_{i}+R_{NT}+C_{NT}$
where $\operatorname*{diag}\{A\}$ indicates the square diagonal matrix with main elements given by the diagonal of $A$, and
$R_{NT}$ $=\sum_{i=1}^{N}(\hat{\beta}_{LS}-\beta)^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}(\hat{\beta}_{LS}-\beta)-2\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}(\hat{\beta}_{LS}-\beta),$
$C_{NT}$ $=2\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})\epsilon_{i}-2\sum_{i=1}^{N}\epsilon_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}(\hat{\beta}_{LS}-\beta).$
Using identical arguments to the ones used for BRS test we have,
$\frac{1}{\sqrt{N}}R_{NT}$ $=O_{p}\left(\frac{M^{2}T}{{N}^{3/2}\gamma^{2}}\right),$
where we explicitly used the fact that $h_{i}$ is a projection matrix, hence $0\leq h_{i,tt}\leq 1$, and
$\sum_{i=1}^{N}\lVert x_{i}^{\prime}\left(h_{i}-\operatorname*{diag}\{h_{i}\}\right)x_{i}\rVert_{2}\leq\sum_{i=1}^{N}\sum_{t=1}^{T}\lVert(x_{it}x_{it}^{\prime})(1-h_{i,tt})\rVert_{2}\leq\sum_{i=1}^{N}\sum_{t=1}^{T}\lVert x_{it}x_{it}\rVert_{2}=O_{p}(NT).$
Here $\lVert A\rVert_{2}=\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}|a_{ij}|^{2}}$ denotes the Frobenius norm of a $k\times k$ matrix $A$. When $M/N\to{m_{0}}>0$ and $MT/\sqrt{N}\gamma^{2}\to c>0$ then
$\frac{1}{\sqrt{N}}R_{NT}=(c)^{2}m_{0}\lambda^{\prime}\left(S^{(G_{2})}S^{-1}WS^{-1}S^{(G_{2})}-W^{(G_{2})}S^{-1}S^{(G_{2})}-S^{(G_{2})}S^{-1}W^{(G_{2})}\right)\lambda+o_{p}(1).$
Similarly,
$\frac{1}{\sqrt{N}}C_{NT}$ $=O_{p}\left(\frac{\sqrt{MT}}{\sqrt{N}\gamma^{2}}\right),$
and if $MT/(\sqrt{N}\gamma^{2})\to c\geq 0$ then $\frac{1}{\sqrt{N}}C_{NT}=o_{p}(1)$. Further,
$\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}\lambda_{i}\xrightarrow{p}\lambda^{\prime}W^{(G_{2})}\lambda\times\lim_{N,T\to\infty}\frac{MT}{\sqrt{N}\gamma^{2}}.$
Finally, applying Theorem 5.1 in BRS, under $H_{1,n}$, we obtain $\hat{V}_{NT}=V_{0}+o_{p}(1)$, and
$\sum_{i=1}^{N}\epsilon_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})\epsilon_{i}\to N(0,V_{0}).$
Putting all of the obtained rates into expression (3) of SC test gives us the non-centrality parter and the statement of the theorem. ∎
Remarks Corollary 1.
The term $C_{NT}$ in all decompositions can be neglected under the conditions of Theorem 1 with $M=\mathbb{M}$. We show the passages to obtain $\delta_{PY}$; the remarks for the other two tests are similar and thus omitted.
Let $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$ where $\lambda_{i}=\sum_{p=2}^{P}\lambda_{p}\mathbb{I}\{g(i)=p\}$, with $\mathbb{I}\{A\}$ being the indicator of event $A$. Then,
$\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}$ $=\frac{1}{\sqrt{N}\gamma^{2}}\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\lambda_{p}=\sum_{p=2}^{P}O_{p}\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)=O_{p}\left(\frac{\mathbb{M}T}{\sqrt{N}\gamma^{2}}\right),$
with
$\mathop{\rm plim}_{N,T\to\infty}\frac{1}{\sqrt{N}\gamma^{2}}\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\lambda_{p}=\lim_{N,T\to\infty}\sum_{p=2}^{P}\lambda_{p}^{\prime}SQ^{(G_{p})}\lambda_{p}\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1).$
Using the same arguments in $R_{NT}$, we obtain,
$\frac{1}{\sqrt{N}}R_{NT}$ $=-\frac{1}{\sqrt{N}\gamma^{2}}\left(\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)\frac{S}{NT}\left(\sum_{q=2}^{P}\lambda_{p}\left(\sum_{i\in G_{q}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)+o_{p}(1)$
$=\sum_{p=2}^{P}\sum_{q=2}^{P}\lambda_{p}^{\prime}S^{(G_{p})}SS^{(G_{q})}\lambda_{q}\left(\frac{M_{p}}{N}\right)\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1)$
Appendix B Formal Statement Fixed $T$
Assumption 3 (Regressors)
  • (a)
    $\mathbb{E}|x_{it,k}|^{4+\delta}<C<\infty$, for some $\delta>0$, and all i=1,,Ni=1,...,N, t=1,,Tt=1,...,T, and k=1,,Kk=1,...,K.
  • (b)
    For any subset $G\subset\{1,...N\}$ such that $G\to\infty$, the probability limits $\Sigma=\mathop{\rm plim}_{N\to\infty}N^{-1}\sum_{i=1}^{N}x_{i}^{\prime}x_{i}$ and $\Sigma^{(G)}=\mathop{\rm plim}_{N\to\infty}(\#G)^{-1}\sum_{i\in G}x_{i}^{\prime}x_{i}$ are positive definite, non-stochastic and finite matrices.
Using the proof of Theorem 1, we obtain the following result:
Corollary 2
Under Assumptions 1, 2, and $H_{1,n}$, as $M,N\to\infty$, the following holds:
$LM\overset{d}{\rightarrow}\begin{cases}\chi_{K}^{2},&\mbox{if }\frac{\sqrt{M}}{\gamma N^{1/4}}\to 0,\\ \chi_{K}^{2}\left(\delta_{LM}^{\prime}\mathcal{V}^{-1}\delta_{LM}\right),&\mbox{if }\frac{\sqrt{M}}{\gamma N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{M}}{\gamma N^{1/4}}\to\infty.\end{cases},$
where $\delta_{LM}$ is a $K\times 1$ vector with $k$-th element
$\delta_{LM,k}=c^{2}\left(\lambda_{2}^{\prime}U_{k}^{(G_{2})}\lambda_{2}+m_{0}\lambda_{2}^{\prime}\left(\Sigma^{(G_{2})}\Sigma^{-1}U_{k}\Sigma^{-1}\Sigma^{(G_{2})}-2\Sigma^{(G_{2})}\Sigma^{-1}U_{k}^{(G_{2})}\right)\lambda_{2}\right),$
with
$U_{k}^{(G)}=\underset{\#G\rightarrow\infty}{\text{plim}}\;(\#G)^{-1}\sum_{i\in G}\sum_{s<t}x_{it,k}x_{it}x_{is,k}x_{is}^{\prime},$
and $U_{k}=U_{k}^{(G_{1}\cup G_{2})}$. Here, $\mathcal{V}$ is a $K\times K$ matrix with $\left(k,l\right)$-th element given by $\mathcal{V}_{k,l}=\underset{N\rightarrow\infty}{\text{plim}}\;N^{-1}\sum_{i=1}^{N}\sum_{s<t}\sigma_{i}^{4}x_{it,k}x_{it,l}x_{is,k}x_{is,l}$.