### v2

# Antonio Raiola

Abstract
Empirical researchers often use slope-homogeneity tests to assess whether slopes can be treated as common across units. A key difficulty is that heterogeneity may be concentrated in a small number of units, so that a failure to reject homogeneity may reflect limited power rather than true homogeneity. We quantify this issue by analyzing the power of standard slope-homogeneity tests under doubly local alternatives — alternatives in which only small groups of units depart from the common slope and the magnitude of the deviations shrinks with sample size. We characterize detectability as a function of panel dimensions, the size of the departing groups, and the rate at which deviations shrink. The results tell the researcher clearly when homogeneity tests are informative and when they will miss small-group heterogeneity. A Monte Carlo study confirms the theory.
JEL classification: C13, C23, C33
Keywords: Panel Data; Detection Boundary; Slope Heterogeneity; Local Power; Group Structures.
1 Introduction
Slope homogeneity is a convenient and widely used restriction in panel data analysis. If slopes are common across units, pooling information over the cross section improves precision and yields estimators with a simple interpretation. If slopes are heterogeneous, common-slope estimands may be difficult to interpret or even misleading (see, e.g., Breitung and Salish, 2021). Empirical researchers therefore use slope-homogeneity tests to assess whether slopes are common and to choose between pooled and heterogeneous specifications. A key difficulty is that heterogeneity may be concentrated in a small number of units, so that non-rejection may reflect limited power against such departures rather than genuine homogeneity. This paper quantifies the local power of standard slope-homogeneity tests against departures confined to small groups. Our results characterize, in terms of panel dimensions and group size, the conditions under which standard tests have power against such localized departures.
The literature on testing slope homogeneity is long-standing. A seminal starting point is Swamy (1970), based on the dispersion of individual-slope estimates around a pooled estimator; refinements such as Pesaran and Yamagata (2008) remain among the most widely used tools for linear panel models. In parallel, several LM-type procedures have been developed in random-coefficient models, where the null of slope homogeneity is equivalent to a zero-variance restriction on the random coefficients. Building on Breusch and Pagan (1979) (see also Hsiao, 2003, Sec. 6.2.2.4), Juhl and Lugovskyy (2014) propose LM statistics for zero variance, and Breitung et al. (2016) further extend the methodology to empirically relevant settings. Other LM tests assess whether the regressors have predictive power for the restricted residuals under the null of slope homogeneity: Su and Chen (2013) in models with interactive fixed effects, and Lu and Su (2017), whose test for the number of latent groups nests slope homogeneity as a special case.
We analyze the power of established test statistics—Pesaran and Yamagata (2008), Su and Chen (2013) [^2] In the setup considered here, the test of Su and Chen (2013) coincides with that of Lu and Su (2017) when specialized to the single-group null; our results therefore cover both simultaneously., and the LM-type test of Breitung et al. (2016)—under a sequence of doubly local alternatives designed to capture small-group heterogeneity. These alternatives are doubly local in the sense that $(i)$ only the slopes of small groups of individuals deviate from the dominant common slope, and $(ii)$ the slope differences across groups shrink at a controlled (Pitman-type) rate. This design characterizes how power depends jointly on the group-specific slope differences and the size of the groups. We develop the analysis for a benchmark setting with two groups (one dominant majority, one heterogeneous remainder), and for extensions with multiple groups, with either a fixed or growing number of groups. The analysis covers both large-$N$, large-$T$ panels and large-$N$, fixed-$T$ panels, where $T$ denotes the time dimension.
Our main results deliver a detection boundary by linking the rate at which slope differences vanish, $\gamma$, to the size of the heterogeneous group(s), $M$. In large-$N$, large-$T$ panels, this boundary is given by
$\gamma\asymp\frac{M^{1/2}T^{1/2}}{N^{1/4}},$
which nests the classical $N^{1/4}\sqrt{T}$ rate—obtained when $M/N\to c>0$—reported in Pesaran and Yamagata (2008), Su and Chen (2013), and Breitung et al. (2016) under analogous Pitman-drift sequences. Importantly, this boundary shows that the tests remain informative even if $M/N\to 0$, provided $MT\to\infty$ and the boundary condition above holds, thereby enabling detection of small-groups alternatives. When $T$ is fixed, detectability requires $M$ to be of order $\sqrt{N}$ or larger; below this threshold, heterogeneity is undetectable. Finally, when the number of heterogeneous groups increases with $N$, the tests retain nontrivial power and detectability depends on the total number of units outside the dominant majority rather than on the number of groups per se.
The results also speak to the literature on latent group structure in panel models, where such departures arise naturally—see, e.g., Bonhomme and Manresa (2015), Bester and Hansen (2016), Su et al. (2016), Lu and Su (2017), and Dzemski and Okui (2024) for identification, estimation, and inference for group membership and group-specific parameters. Our contribution is complementary, as we characterize when standard slope-homogeneity tests can detect departures consistent with latent group structure. This perspective also connects to recent work on inference when latent groups are not well separated; see Akgun and Okui (2025).
The remainder of the paper is organized as follows. Section 2 introduces the model and notation. Section 3 reviews the test statistics. Section 4 derives the local-power results under doubly local alternatives, and Section 5 reports a Monte Carlo study. Proofs are relegated to the appendix.
2 Preliminaries
Consider the linear panel data model
$y_{i}=x_{i}^{\prime}\beta_{g(i)}+\epsilon_{i},\quad i=1,\ldots,N,$ (1)
where $y_{i}=(y_{i1},\ldots,y_{iT})^{\prime}$ is the $T$-dimensional vector of dependent variables for unit $i$; $x_{i}=(x_{i1},\ldots,x_{iT})^{\prime}$ is the $T\times K$ design matrix of regressors for unit $i$; and $\epsilon_{i}=(\epsilon_{i1},\ldots,\epsilon_{iT})^{\prime}$ is the corresponding vector of idiosyncratic errors. The slope parameter, $\beta_{g(i)}$, is group-specific, with unknown group membership determined by the mapping $g:\mathcal{N}\to\{G_{1},\ldots,G_{P}\}$, where $\mathcal{N}=\{1,\ldots,N\}$. The function $g$ assigns unit $i$ to one of the $P$ groups partitioning the sample; that is, $G_{p}\cap G_{q}=\emptyset$ for all $p\neq q$, and $\bigcup_{p=1}^{P}G_{p}=\mathcal{N}$.
We are interested in testing the presence of group heterogeneity in the slopes by means of existing slope homogeneity tests. The null hypothesis of identical slopes is
$H_{0}:\beta_{g(i)}=\beta,\text{ for all }i\in\mathcal{N},$
versus the alternative that there exists a group of individuals with different slopes,
$\G.H_{1}:\exists\ G\subset\{1,...,N\}:\;G\neq\emptyset\;\text{and}\;\beta_{g(i)}\neq\beta_{g(j)}\;\text{for}\;i\in G\;\text{and}\;j\in\mathcal{N}\backslash$ G.
As discussed in the introduction, a variety of well-established tests for slope homogeneity are available in the literature. To analyze their asymptotic behavior within a unified framework, we adopt a common set of assumptions sufficiently broad to capture their essential features. Throughout, $C$ denotes a generic constant that may vary across expressions, and #G\#G denotes the cardinality of a generic set $G\subseteq\mathcal{N}$.
Assumption 1 (Errors)
$(a)$ $\epsilon_{it}|x_{i}\sim IID(0,\sigma^{2}_{i})$, where $\max_{i}\sigma^{2}_{i}<C<\infty$, $\min_{i}\sigma^{2}_{i}>0$, with $\epsilon_{it}$ and $\epsilon_{js}$ indipendently distributed for $i\neq j$ and $t\neq s$; $(b)$ $\mathbb{E}\left[\epsilon_{it}^{6}|x_{i}\right]<C<\infty$.
Assumption 2 (Regressors)
For all $i\in\mathcal{N}$: $(a)$ $\mathbb{E}|x_{it,k}|^{6}<C<\infty$, for all t=1,,Tt=1,...,T and k=1,,Kk=1,...,K; $(b)$ the probability limit $S_{i}=\mathop{\rm plim}_{T\to\infty}T^{-1}x_{i}^{\prime}x_{i}$, is a positive definite, non-stochastic and finite matrix; (c) for any subset $G\subseteq\mathcal{N}$ such that $G\to\infty$, the probability limit $S^{(G)}=\mathop{\rm plim}_{N,T\to\infty}(\#GT)^{-1}\sum_{i\in G}x_{i}^{\prime}x_{i}$ is a positive definite, non-stochastic and finite matrix.
In particular, taking $G=\mathcal{N}$, $S=S^{(\mathcal{N})}$ is a positive definite non-stochastic matrix.
Assumption 1(a), corresponding to Assumption 1(i)–(ii) in Pesaran and Yamagata (2008), rules out serial correlation, and cross-sectional dependence. Together with model (1), which excludes endogenous components in the error term, they provide a simplified framework for presenting our main results in a unified manner. We emphasize, however, that the tests considered here have been developed or extended to accommodate more general error structures (see original papers as well as Blomquist and Westerlund 2013; Ando and Bai 2015). In all such cases, the main conclusions of our analysis remain valid.
We focus here on three representative and widely used slope homogeneity tests: (i) the dispersion statistic $\Delta$ from Pesaran and Yamagata (2008) (hereafter, PY); (ii) the residual-based LM test $\mathcal{J}$ from Su and Chen (2013) (SC); and (iii) the LM-type variance test $LM$ from Breitung et al. (2016) (BRS). The next section provides the necessary details for each of the test statistics.
3 Test Statistics
The $\Delta$ test of PY refines the Swamy (1970) test, originally designed for testing slope homogeneity in large $T$, fixed $N$, scenarios. Both tests compare individual slope estimates, $\hat{\beta}_{i}=(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}y_{i}$, with the weighted least squares (WLS) estimator $\hat{\beta}_{WLS}(\bar{\sigma}_{i}^{2})=\bigl(\sum_{i=1}^{N}x_{i}^{\prime}x_{i}/\bar{\sigma}_{i}^{2}\bigr)^{-1}\bigl(\sum_{i=1}^{N}x_{i}^{\prime}y_{i}/\bar{\sigma}_{i}^{2}\bigr)$, using some estimator $\bar{\sigma}^{2}_{i}$ of $\sigma^{2}_{i}$. The original Swamy statistic is $\mathcal{S}_{SW}=\mathcal{S}(\hat{\sigma}^{2}_{i})$, where
$\mathcal{S}(\bar{\sigma}^{2}_{i})=\sum_{i=1}^{N}\left(\hat{\beta}_{i}-\hat{\beta}_{WLS}(\hat{\sigma}^{2}_{i})\right)^{\prime}\left(\frac{x_{i}^{\prime}x_{i}}{\bar{\sigma}^{2}_{i}}\right)\left(\hat{\beta}_{i}-\hat{\beta}_{WLS}(\hat{\sigma}^{2}_{i})\right),$
and $\hat{\sigma}^{2}_{i}=(T-K-1)^{-1}(y_{i}-x_{i}\hat{\beta}_{i})^{\prime}(y_{i}-x_{i}\hat{\beta}_{i})$. This statistic has a $\chi^{2}$ null distribution with K(N1)K(N-1) degrees of freedom when $T\to\infty$ and $N$ is fixed. For the large-$N$, large-$T$ framework, PY propose to replace $\hat{\sigma}^{2}_{i}$ with
$\widetilde{\sigma}^{2}_{i}=\frac{\hat{\epsilon}_{i}^{\prime}\hat{\epsilon}_{i}}{T-K-1},$
where $\hat{\epsilon}_{i}=(\hat{\epsilon}_{i1},...,\hat{\epsilon}_{iT})^{\prime}$, $\hat{\epsilon}_{it}=y_{it}-x_{it}^{\prime}\hat{\beta}_{LS}$, and $\hat{\beta}_{LS}=(\sum_{i=1}^{N}x_{i}^{\prime}x_{i})^{-1}\sum_{i=1}^{N}x_{i}^{\prime}y_{i}$ - the pooled least square estimator. The resulting dispersion statistic, $\mathcal{S}_{PY}=\mathcal{S}(\tilde{\sigma}^{2}_{i})$, after centering and standardization, converges to the standard normal distribution under $H_{0}$,
$\Delta=\frac{\mathcal{S}_{PY}-NK}{\sqrt{2NK}}\overset{d}{\rightarrow}\mathcal{N}(0,1),\quad\text{ as }\quad N,T\to\infty.$ (2)
The residual-based $\mathcal{J}$ test. An alternative approach is the LM test of SC, [^3] With additive fixed effects (or none, as in the present case), $\mathcal{J}$ is numerically equivalent to the test of Lu and Su (2017) when applied to a single-group case under the null. based on the idea that, under $H_{0}$ and with predetermined regressors, $x_{it}$ should not have explanatory power for the residuals in (1). This motivates a test of $H_{0}^{*}:\phi_{i}=0\;\text{for all }i=1,\ldots,N$ in the auxiliary regression
$\hat{\epsilon}_{it}=\phi_{i}^{\prime}x_{it}+\eta_{it},\qquad\text{with}\quad i=1,\ldots\,N,t=1,\ldots,T,$
where $\eta_{it}$ is a mean-zero error term.
Let $h_{i}=x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}$ be the projection matrix with diagonal elements $h_{i,tt}$, and define $\hat{b}_{it}=\hat{S}_{i}^{-1/2}x_{it}$, where $\hat{S}_{i}=T^{-1}x_{i}^{\prime}x_{i}$. The SC test statistic is given by the following centered and standardized $LM$ test statistic, which has a limiting normal distribution under $H_{0}^{*}$:
$\mathcal{J}=\frac{N^{-1/2}LM_{SC}-\hat{B}_{NT}}{\sqrt{\hat{V}_{NT}}}\overset{d}{\rightarrow}\mathcal{N}(0,1),\quad N,T\to\infty,$ (3)
where
$LM_{SC}=\sum_{i=1}^{N}\hat{\epsilon}_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\hat{\epsilon}_{i},$
and $\hat{B}_{NT}=N^{-1/2}\sum_{i=1}^{N}\sum_{t=1}^{T}\hat{\epsilon}_{it}^{2}h_{i,tt}$ and $\hat{V}_{NT}=4T^{-2}N^{-1}\sum_{i=1}^{N}\sum_{t=2}^{T}\left[\hat{\epsilon}_{it}\hat{b}_{it}^{\prime}\sum_{s=1}^{t-1}\hat{b}_{is}\hat{\epsilon}_{is}\right]^{2}$ estimate the mean and variance of $LM_{SC}$ under the null, respectively.
The LM variance test. A third approach reformulates the null hypothesis within the random coefficients model. Specifically, assuming $\beta_{i}=\beta+v_{i}$ with $\mathbb{E}[v_{i}v_{i}^{\prime}|x_{i}]=\text{diag}(\sigma^{2}_{v,1},\dots,\sigma^{2}_{v,K})$, slope homogeneity is equivalent to $H_{0}^{**}:\sigma^{2}_{v,k}=0$, for all $k=1,\dots,K$, which reduces the N(K1)N(K-1) linear restrictions tested by $\Delta$ and $\mathcal{J}$ to $K$.
Building on this idea, BRS develop an LM test valid for panels with large $N$ and either fixed $T$ or large $T$. The statistic is based on the score vector $\hat{s}=(\hat{s}_{1},\ldots,\hat{s}_{K})^{\prime}$, where
$\hat{s}_{k}=\sum_{i=1}^{N}\sum_{t=2}^{T}\sum_{s=1}^{t-1}\hat{\epsilon}_{it}\hat{\epsilon}_{is}x_{it,k}x_{is,k},\;\;k=1,...,K.$
The test statistic is then given by
$LM=\hat{s}^{\prime}\hat{V}^{-1}\hat{s}$, (4)
where $\hat{V}$ has $(k,l)$-th element
$\hat{V}_{kl}=\sum_{i=1}^{N}\sum_{t=2}^{T}\sum_{s=1}^{t-1}\sum_{p=1}^{t-1}\hat{\epsilon}_{it}^{2}x_{it,k}x_{it,l}\,\hat{\epsilon}_{is}\hat{\epsilon}_{ip}\,x_{is,k}x_{ip,l}\quad\text{and}\quad l,k=1,...,K.$
Under $H_{0}^{**}$ and $N\to\infty$, with $T$ either fixed or $T\to\infty$, the test has a $\chi^{2}_{K}$ limiting distribution.
The three tests, $\Delta$, $\mathcal{J}$, and $LM$, are consistent for $H_{0}$ and have well-established local power properties. However, none has been explicitly studied for local power against small-group heterogeneity, which is the focus of the next section.
4 Asymptotic Local Power Analysis
We study the local power of $\Delta$, $\mathcal{J}$, and $LM$ tests under grouped alternatives, focusing on (i) how large slope differences must be to be detected and (ii) how large the groups must be for detection. To this end, we introduce a sequence of doubly local alternatives that extends the classical Pitman framework and integrates these two aspects. We first present the results for the case where the sample is partitioned into two groups, $\{G_{1},G_{2}\}$, with $N,T\to\infty$. Group $G_{1}$ is treated as the “dominant” group and $G_{2}$ as the “alternative” one, containing the remaining individuals. We then extend these results to panels with $N\to\infty$ and fixed $T$, as well as to settings with more than two groups.
Formally, define
$H_{1,n}:\beta_{g(i)}=\begin{cases}\beta,&\text{if }g(i)=G_{1},\\ \beta+\lambda_{2}/\gamma,&\text{if }g(i)=G_{2},\end{cases}$
where $\lambda_{2}\neq 0$ is a $K\times 1$ vector, and $\{G_{1},G_{2}\}$ form a partition of $\mathcal{N}$. The number of individuals in each group is
$G_{2}=M,\qquad\#G_{1}=N-M,$
where $M\to\infty$ as $N\to\infty$ and $\gamma=\gamma(N,T,M)$ diverges with $N$, $T$, and $M$.
Allowing the scaling factor $\gamma$ to depend on $M$ as well as on $N$ and $T$ reflects that our proposed local alternatives control for both the slope difference and the size of the heterogeneous group. This extension is essential: the relative magnitude of $M$ with respect to $N$ and $T$ determines when slope homogeneity tests can detect group-based heterogeneity. Theorem 1 formalizes this statement.
Theorem 1
Under Assumptions 1, 2, and $H_{1,n}$, as $M,N,T\to\infty$ and $M/N\to{m_{0}}\in[0,1)$, the following holds:
(PY test)
Additionally assume $\mathbb{E}\left[\epsilon_{it}^{9}|x_{i}\right]<C<\infty$ and $\sqrt{N}/T^{2}\to 0$, then,
$\Delta\overset{d}{\rightarrow}\begin{cases}\mathcal{N}(0,1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to 0,\\ \mathcal{N}(\delta_{\Delta},1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to\infty.\end{cases},$
where
${\delta}_{\Delta}=\frac{c^{2}}{\sqrt{2K}}\lambda_{2}^{\prime}\left(Q^{(G_{2})}-m_{0}Q^{(G_{2})}Q^{-1}Q^{(G_{2})}\right)\lambda_{2},$
with $Q^{(G)}=\mathop{\rm plim}\limits_{\#G,T\to\infty}(\#GT)^{-1}\sum_{i\in G}x_{i}^{\prime}x_{i}/\sigma^{2}_{i}$, and $Q=Q^{(\mathcal{N})}$.
(SC test)
$\mathcal{J}\overset{d}{\rightarrow}\begin{cases}\mathcal{N}(0,1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to 0,\\ \mathcal{N}(\delta_{\mathcal{J}},1),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to\infty.\end{cases},$
where
${\delta}_{\mathcal{J}}=\frac{c^{2}}{\sqrt{V_{0}}}\lambda_{2}^{\prime}\left(W^{(G_{2})}+m_{0}(S^{(G_{2})}S^{-1}WS^{-1}S^{(G_{2})}-2W^{(G_{2})}S^{-1}S^{(G_{2})})\right)\lambda_{2},$
with $W^{(G)}=\mathop{\rm plim}\limits_{\#G,T\to\infty}(\#GT)^{-1}\sum_{i\in G}\sum_{t=1}^{T}{x_{it}x_{it}^{\prime}}(1-h_{i,tt})$, and $W=W^{(\mathcal{N})}$. Here $V_{0}=\lim_{N,T\to\infty}4T^{-2}N^{-1}\sum_{i=1}^{N}\sum_{t=2}^{T}\mathbb{E}\left[\left(\epsilon_{it}b_{it}^{\prime}\sum_{s=1}^{t-1}b_{is}\epsilon_{is}\right)^{2}\right]$, with $b_{it}=S^{-1/2}_{i}x_{it}$.
(BRS test)
$LM\overset{d}{\rightarrow}\begin{cases}\chi_{K}^{2},&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to 0,\\ \chi_{K}^{2}\left(\delta_{LM}^{\prime}\Psi^{-1}\delta_{LM}\right),&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{MT}}{\gamma\,N^{1/4}}\to\infty.\end{cases},$
where $\delta_{LM}$ is a $K\times 1$ vector, with $k$-th element
$\delta_{LM,k}=c^{2}\left(\lambda_{2}^{\prime}\Omega_{k}^{(G_{2})}\lambda_{2}+m_{0}\lambda_{2}^{\prime}\left(S^{(G_{2})}S^{-1}\Omega_{k}S^{-1}S^{(G_{2})}-2S^{(G_{2})}S^{-1}\Omega_{k}^{(G_{2})}\right)\lambda_{2}\right)$
with $\Omega_{k}^{(G)}=\mathop{\rm plim}_{\#G,T\rightarrow\infty}\;\frac{1}{\#GT^{2}}\sum_{i\in G}\sum_{s<t}x_{it,k}x_{it}x_{is,k}x_{is}^{\prime}$, and $\Omega_{k}=\Omega_{k}^{(\mathcal{N})}$. Here, $\Psi$} is a $K\times K$ matrix with ${\Psi}_{k,l}=\underset{N,T\rightarrow\infty}{\text{plim}}\;(NT^{2})^{-1}\sum_{i=1}^{N}\sum_{s<t}\sigma^{4}_{i}x_{it,k}x_{it,l}x_{is,k}x_{is,l}$.
Theorem 1 highlights a trade-off between the rate at which slope differences vanish and the size of the alternative group. The critical rate,
$\gamma\asymp\frac{M^{1/2}T^{1/2}}{N^{1/4}},\qquad M,N,T\to\infty,$ (5)
marks the boundary between trivial and nontrivial power. If $\gamma$ grows faster than this rate, all three tests have trivial power. If it grows at or below this rate, they have power to detect the grouped alternative.
Two cases deserve attention. If the fraction of alternative units remains bounded away from zero ($m_{0}>0$, so $M\asymp N$), the rate (5) reduces to the classical order $O(\sqrt{T}N^{1/4})$, consistent with the local power results of PY, SC, and BRS. When the alternative group is small ($m_{0}=0$), the rate is slower than $O(\sqrt{T}N^{1/4})$, yet the tests retain nontrivial power as long as $M$ diverges with $N$ — even as the fraction of units with distinct slopes vanishes. This contrasts with the prevailing assumption in the grouped-patterns literature, which requires heterogeneous slopes to affect a non-negligible share of the sample (e.g., Bonhomme and Manresa, 2015; Su et al., 2016; Lu and Su, 2017).
On the noncentrality parameters.
Theorem 1 shows that all three tests share the same asymptotic detection boundary (5). Their local power need not coincide, however, since the noncentrality parameters differ. A general comparison is intractable, as the parameters depend on both model and design in complex ways. We therefore consider the benchmark case of a single regressor, $K=1$, to build intuition.
Let $x_{it}$ be scalar and i.i.d. across $i$ and $t$, with $\mathbb{E}\left[x_{it}\right]=0$, $\mathbb{E}\left[x_{it}^{2}\right]=\sigma_{i,x}^{2}$, and uniformly bounded fourth moments. The regressor may thus exhibit unit-specific variance, constant over time. For any subset $G\subseteq\mathcal{N}$, define
$\mu_{2,x}^{(G)}=\lim_{\#G\to\infty}\frac{1}{\#G}\sum_{i\in G}\sigma_{i,x}^{2},\quad\mu_{4,x}^{(G)}=\lim_{\#G\to\infty}\frac{1}{\#G}\sum_{i\in G}\bigl(\sigma_{i,x}^{2}\bigr)^{2},\quad q_{x}^{(G)}=\lim_{\#G\to\infty}\frac{1}{\#G}\sum_{i\in G}\frac{\sigma_{i,x}^{2}}{\sigma_{i}^{2}},$
with $\mu_{2,x}=\mu_{2,x}^{(\mathcal{N})}$, $\mu_{4,x}=\mu_{4,x}^{(\mathcal{N})}$, and $q_{x}=q_{x}^{(\mathcal{N})}$. Now suppose $m_{0}=0$, so $M/N\to 0$ and the heterogeneous subset is asymptotically negligible. The second term in each noncentrality parameter vanishes, and Theorem 1 reduces to
$\delta_{\Delta}=\frac{c^{2}\lambda_{2}^{2}}{\sqrt{2}}\,q_{x}^{(G_{2})},\qquad\delta_{\mathcal{J}}=\frac{c^{2}\lambda_{2}^{2}}{\sqrt{V_{0}}}\,\mu_{2,x}^{(G_{2})},$
and
$\xi_{LM}=\delta_{LM}^{\prime}\Psi^{-1}\delta_{LM}=\frac{c^{4}\lambda_{2}^{4}}{4\Psi}\,\left(\mu_{4,x}^{(G_{2})}\right)^{2},$
where $V_{0}=2\lim_{N\to\infty}N^{-1}\sum_{i=1}^{N}\sigma_{i}^{4}$ and $\Psi=\lim_{N\to\infty}(2N)^{-1}\sum_{i=1}^{N}\sigma_{i}^{4}(\sigma_{i,x}^{2})^{2}$. Even in the scalar case, PY and SC depend on different weighted second moments, while BRS depends on fourth-order cross-sectional moments of the regressor.
To compare PY and SC directly, note that
$\frac{\delta_{\Delta}}{\delta_{\mathcal{J}}}=\lim_{N\to\infty}\mathrm{RMS}(\sigma^{2}_{i})\sum_{i\in G_{2}}\frac{\omega_{x,i}}{\sigma^{2}_{i}}\;$
where $\omega_{x,i}=\sigma^{2}_{x,i}/\sum_{i\in G_{2}}\sigma^{2}_{x,i}$, and $\mathrm{RMS}(\sigma^{2}_{i})=\sqrt{N^{-1}\sum_{i=1}^{n}\sigma^{4}_{i}}$. It follows that,
$\lim_{N\to\infty}\frac{\mathrm{RMS}(\sigma^{2}_{i})}{\max_{i\in G_{2}}\sigma^{2}_{i}}\leq\frac{\delta_{\Delta}}{\delta_{\mathcal{J}}}\leq\lim_{N\to\infty}\frac{\mathrm{RMS}(\sigma^{2}_{i})}{\min_{i\in G_{2}}\sigma^{2}_{i}}.$
When $G_{2}$ concentrates on units with small $\sigma^{2}_{i}$, the $\Delta$ test dominates; otherwise the $\mathcal{J}$ test has higher local power. Under homoskedasticity ($\sigma_{i}^{2}=\sigma^{2}$ for all $i$), $\delta_{\Delta}=\delta_{\mathcal{J}}$ and the two tests are locally equivalent.
Comparing PY and SC with BRS is more complex. Under homoskedasticity and constant regressor variance ($\sigma_{i,x}^{2}=\sigma_{x}^{2}$ for all $i$), all three procedures are locally equivalent: $\xi_{LM}/\delta_{\Delta}^{2}=\xi_{LM}/\delta_{\mathcal{J}}^{2}=1$. If errors are homoskedastic and $\{\sigma_{x,i}^{2}\}$ unrestricted with $\mu_{2,x}^{(G_{2})}=\mu_{2,x}$, and $\mu_{4,x}^{(G_{2})}=\mu_{4,x}$, then
$\frac{\xi_{LM}}{\delta_{\Delta}^{2}}=\frac{\mu_{4,x}}{\mu_{2,x}^{2}}\geq 1.$
Hence, BRS has a larger local noncentrality parameter than PY and SC, with the ratio $\xi_{LM}/\delta_{\Delta}^{2}$ governed by the cross-sectional dispersion in regressor variance.
A complementary comparison is obtained when regressor variance is constant across units but errors are heteroskedastic. Then, $\delta_{\mathcal{J}}^{2}=\xi_{LM}$, so SC and BRS tests are locally equivalent. If $G_{2}$ is also asymptotically representative with respect to error variances — that is $\lim_{M\to\infty}M^{-1}\sum_{i\in G_{2}}\sigma_{i}^{-2}=\lim_{N\to\infty}N^{-1}\sum_{i=1}^{n}\sigma_{i}^{-2}$ — then
$\frac{\xi_{LM}}{\delta_{\Delta}^{2}}=\left(\frac{\delta_{\mathcal{J}}}{\delta_{\Delta}}\right)^{2}=\lim_{N\to\infty}\frac{1}{\left(N^{-1}\sum_{i=1}^{N}\sigma_{i}^{4}\right)\left(N^{-1}\sum_{i=1}^{N}\sigma_{i}^{-2}\right)^{2}}\leq 1,$
with equality under homoskedasticity. When regressor variance is constant, error heteroskedasticity thus favors PY over SC and BRS
When $m_{0}>0$ (i.e., $M\asymp N$), the comparison is more complex as the second term in each expression of noncentrality parameters is no longer asymptotically negligible. However, the analysis of the same scalar benchmark points in the same direction. In particular, under homoskedasticity, $\delta_{\Delta}=\delta_{\mathcal{J}}$; if, in addition, the regressor variance is constant, all three tests are locally equivalent; if, instead, regressor variance is unrestricted and $\mu_{2,x}^{(G_{2})}=\mu_{2,x}$, and $\mu_{4,x}^{(G_{2})}=\mu_{4,x}$,
$\frac{\xi_{LM}}{\delta_{\Delta}^{2}}=\frac{\mu_{4,x}}{\mu_{2,x}^{2}}\geq 1.$
Outside these benchmarks, the comparison depends on the joint distribution of $\{\sigma^{2}_{i}\}$ and $\{\sigma^{2}_{x,i}\}$ across the full sample and the remainder group $G_{2}$.
4.1 Panels with $N\to\infty$ and $T$-fixed
When $T$ is fixed and $N$ grows, the LM-type test of BRS is the only one available. The restrictions on the regressors can be relaxed, as in Assumption 3 in the Appendix, which also contains the corresponding version of Theorem 1. In this scenario, the boundary between trivial and nontrivial power is governed by the rate
$\gamma\asymp\frac{\sqrt{M}}{N^{1/4}},\qquad M,N\to\infty,$ (6)
the analogue of (5) in this setting.
When the alternative group is proportional to the total cross-section ($m_{0}>0$, or $M\!\asymp\!N$), the rate in (6) reduces to $\gamma\!\asymp\!N^{1/4}$, matching the local-power behavior established in BRS. More interestingly, this rate marks the boundary at which the test begins to exhibit power. If $M$ diverges at least as fast as $N^{1/2}$, the $LM$ test retains nontrivial local power even when $M/N\!\to\!0$. Essentially, the minimal detectable group size is of order $N^{1/2}$: below this threshold, grouped heterogeneity is undetectable.
4.2 Panels with increasing number of groups
The main message of our analysis remain valid even if the number of groups is allowed to grow with $N$. Consider
$H_{2,n}:\beta_{g(i)}=\begin{cases}\beta,&\mbox{if }g(i)=G_{1},\\ \beta+\lambda_{p}/\gamma,&\mbox{if }g(i)=G_{p}\mbox{ and }p=2,...,P,\end{cases}$
where $\{G_{1},...,G_{P}\}$ forms a partition of $\mathcal{N}$, with $\lambda_{p}\neq\lambda_{q}$ for any $p,q\in\{1,...,P\}$ such that $p\neq q$. We let $M_{p}=\#G_{p}$ and $\mathbb{M}=\sum_{p=2}^{P}M_{p}$.
Analyzing the local power when $P\to\infty$ as $N\to\infty$ is complicated because the size of each group $\{G_{2},\dots,G_{P}\}$ can also depend on $N$. Both the number of groups and the number of individuals per group determine the expansion rates under which the test statistics exhibit nontrivial power.
Corollary 1
Let Assumptions 1, 2, and the sequence of local alternatives $H_{2,n}$ hold. If $\mathbb{M}\to\infty$ as $N\to\infty$, then the result of Theorem 1 holds true by replacing $M$ with $\mathbb{M}$, and the non-centrality parameters, $\delta_{\Delta}$, $\delta_{\mathcal{J}}$, and $\delta_{LM,k}$ with,
$\lim_{N,T\to\infty}\frac{1}{\sqrt{2K}}\left(\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}Q^{(G_{p})}\lambda_{p}-\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}\left(Q^{(G_{p})}Q^{-1}Q^{(G_{q})}\right)\lambda_{q}\right),$
$\lim_{N,T\to\infty}\frac{1}{\sqrt{V_{0}}}\left(\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}W^{(G_{p})}\lambda_{p}+\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}\left[S^{(G_{p})}S^{-1}WS^{-1}S^{(G_{q})}-2W^{(G_{p})}S^{-1}S^{(G_{q})}\right]\lambda_{q}\right)$
and,
$\lim_{N,T\to\infty}\sum_{p=2}^{P}c_{p}^{2}\lambda_{p}^{\prime}U_{k}^{(G_{p})}\lambda_{p}+\sum_{p,q=2}^{P}m_{p}c_{q}^{2}\lambda_{p}^{\prime}\left(S^{(G_{p})}S^{-1}\Omega_{k}S^{-1}S^{(G_{q})}-2S^{(G_{p})}S^{-1}\Omega_{k}^{(G_{q})}\right)\lambda_{q},$
respectively, where for all p=2,,Pp=2,...,P, $M_{p}/N\to m_{p}\in[0,1)$ and $M_{p}T/(\gamma^{2}\sqrt{N})\to c_{p}^{2}\in[0,c^{2}]$, with $c=\mathbb{M}T/(\sqrt{N}\gamma^{2})$.
Corollary 1 generalizes Theorem 1. If $P$ is fixed and $M^{\ast}=\max_{2\leq p\leq P}M_{p}\to\infty$ as $N\to\infty$, replacing $M$ with $M^{\ast}$ yields Theorem 1. If $P\to\infty$ while $\sup_{2\leq p\leq P}M_{p}=\kappa<\infty$, the expansion rate satisfies $PT/(\sqrt{N}\gamma^{2})\to c>0$ as $N,T\to\infty$. More generally, when both $P\to\infty$ and $M_{p}\to\infty$ for at least one $p$, the expansion rate is determined by the total “volume” of individuals in $\{G_{2},\dots,G_{P}\}$, denoted as $\mathbb{M}$.
5 Monte Carlo Study
We evaluate the finite-sample performance of the tests by comparing their power and assessing the theoretical results presented in the main text. We consider the three test statistics ($\Delta$, $\mathcal{J}$, and $LM$) under a data-generating process adapted from PY to accommodate grouped slope patterns; i.e., for each $i=1,\ldots,N$ and $t=1,\ldots,T$:
$y_{it}$ $=x_{it}^{\prime}\beta_{g(i)}+\alpha_{i}+\epsilon_{it},$
$x_{it,k}$ $=\rho_{i,k}x_{it-1,k}+\sqrt{1-\rho_{i,k}^{2}}\,v_{it,k},\qquad k=1,\ldots,K,$
where $x_{it}=(x_{it,1},\ldots,x_{it,K})^{\prime}$, $\epsilon_{it}\overset{iid}{\sim}\mathcal{N}(0,1)$, $v_{it}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{ix}^{2})$, $\alpha_{i}\overset{iid}{\sim}\mathcal{N}(1,1)$, and $\rho_{i}\overset{iid}{\sim}\mathcal{U}(0.05,0.95)$, with $\sigma_{ix}^{2}\sim\chi^{2}_{1}$. The first 50 observations of $x_{it}$ ($t=-49,\ldots,0$) are discarded to mitigate the effect of initial values. Group-specific slopes, $\beta_{g(i)}=(\beta_{g(i),1},...,\beta_{g(i),K})^{\prime}$, are generated as:
$\beta_{g(i),k}=\begin{cases}1,&\text{if }g(i)=G_{1},\\ 1+\lambda(1+e_{p}),&\text{if }g(i)=G_{p},\end{cases}$
for $k=1,\ldots,K$, where $p=2,\ldots,P$ and $e_{p}\overset{iid}{\sim}\mathcal{U}(-h,h)$. The dominant group $G_{1}$ contains $N-\sum_{p=2}^{P}M_{p}$ individuals, while the remaining P1P-1 groups $\{G_{2},\ldots,G_{P}\}$ represent alternative groups of size $M_{p}$. The parameters $\lambda$ and $h$ control, respectively, the degree of heterogeneity relative to the dominant group and the dispersion across alternative groups.
We consider two simulation designs to examine how the size of the alternative group and the magnitude of slope differences affect power:
  • (1)
    Two-group design: $P=2$, $h=0$, varying the size of the alternative group $M_{2}$ for $\lambda\in\{0.1,0.2,0.3\}$.
  • (2)
    Multiple-group design: $P\in\{3,5,10\}$, where alternative groups size are randomly assigned subject to $\sum_{p=2}^{P}M_{p}=\mathbb{M}$, with varying $\mathbb{M}$. Here, $\lambda=0.2$ and $h=0.2$, focusing on the effect of the overall mass of alternative groups $\mathbb{M}$.
Each design is run at $(N,T)=(100,10)$ and $(100,100)$, corresponding to short and long panels, and at $K=1$ and $K=4$, corresponding to single and multiple regression.
In both designs, the $\mathcal{J}$ test is severely biased when $T$ is small, especially for $K=4$, and its rejection rate far exceeds that of the other two tests. To compare the three tests on equal footing, we use size-adjusted critical values obtained by simulating the null distribution of each test with $R_{2}=10000$ replications.
The data-generating process includes individual effects, which are removed through appropriate demeaning procedures for each test statistic. For the $LM$ test, we employ orthogonal deviations as in Arellano and Bover (1995), following Remark 5.1 in BRS. For the $\Delta$ and $\mathcal{J}$ tests, we apply the standard within transformation, pre-multiplying $y_{i}$ and $x_{i}$ by $M_{0}=I_{T}-T^{-1}\mathbf{i}_{T}\mathbf{i}_{T}^{\prime}$, where $\mathbf{i}_{T}$ is a $T\times 1$ vector of ones, as suggested in PY and Lu and Su (2017).
Two Groups: $K=1$, $N=100$, $T=10$
Two Groups: $K=1$, $N=100$, $T=100$
Two Groups: $K=4$, $N=100$, $T=10$
Two Groups: $K=4$, $N=100$, $T=100$
Figure 1: Power curves of slope homogeneity tests with two groups.
$P$ Groups: $K=1$, $N=100$, $T=10$
$P$ Groups: $K=1$, $N=100$, $T=100$
$P$ Groups: $K=4$, $N=100$, $T=10$
$P$ Groups: $K=4$, $N=100$, $T=100$
Figure 2: Power curves of slope homogeneity tests with two groups.
All simulations report rejection frequencies at the 5% significance level based on 1,000 Monte Carlo replications. Results for the two-group design are shown in Figure 1, and those for the multiple-group design in Figure 2.
The two-group results closely mirror our theoretical findings in Theorem 1. Rejection rates increase monotonically with the alternative group size $M$ and with the slope deviation parameter $\lambda$. All tests tend to have a power close to trivial when $M$ is smaller than $\sqrt{N}$, but the power increases significantly after this threshold and as $T$ grows.
In the multiple-group experiment, the decisive role is played by the number of individuals in the alternative groups, $\mathbb{M}$. Tests exhibit substantial power when $\mathbb{M}$ grows at least proportionally to $\sqrt{N}$, consistent with Corollary 1. Power declines slightly as $P$ increases, due to greater dispersion in group-specific slopes, but this effect is secondary once $\mathbb{M}$ or $T$ is sufficiently large.
In most cases, all three tests show comparable power. The $J$ and $LM$ tests, however, tend to gain power somewhat more rapidly than the $\Delta$ test, especially in the short-panel designs with $T=10$ and $K=4$. This divergence is much less visible once $T$ increases. When $T=100$, the three procedures behave in a way that is broadly consistent with the asymptotic comparison in Section 4.
6 Conclusions
We derive the detection boundary of standard slope-homogeneity tests under small-group heterogeneity. The power of the tests is evaluated under a sequence of doubly local alternatives, linking the degree of heterogeneity across groups and each group size to its impact on test performance. Our analysis shows that slope-homogeneity tests are effective in detecting such structures, including sparse groups whose cardinality is negligible relative to the cross-sectional size. Monte Carlo simulations support similar conclusions in finite samples.
Appendix
Appendix A Proofs
We derive the asymptotic distribution of the three type of tests in Theorems 1 under $H_{1,n}$, in the large panel setting, where $N,T\to\infty$. The rates stated in Corollaries 1 and 2 (at the end of the Appendix) follow immediately by setting $M=\mathbb{M}$ and $T=1$, respectively. Additional remarks on the derivation of the noncentrality parameters in Corollary 1 are provided at the end of this section.
The proof is based on the decomposition of each test into a term leading the asymptotic distribution of the test, a term that drives the power of the test, and remainder terms, denoted as $R_{NT}$ and $C_{NT}$, that are asymptotically negligible under the conditions of the statements.
Under $H_{1,n}$, we can rewrite model (1) as
$y_{it}=x_{it}^{\prime}\beta_{i}+\epsilon_{it}=x_{it}^{\prime}\beta+x_{it}^{\prime}\lambda_{i}+\epsilon_{it}=x_{it}^{\prime}\beta+u_{it},$
where $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$, $\lambda_{i}=0$ if $i\in G_{1}$ and $\lambda_{i}=\lambda/\gamma$ if $i\in G_{2}$.
Lemma 1
Under $H_{1,n}$ and Assumptions 1, 2,
$\hat{\beta}_{LS}-\beta=O_{p}\left(\frac{M}{N\gamma}+\frac{1}{\sqrt{NT}}\right),\,\mbox{as }N,T\to\infty.$
Proof of Lemma 1.
$\hat{\beta}_{LS}-\beta$ $=$ $\left(\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\right)^{-1}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}\epsilon_{it}+\left(\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\right)^{-1}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\lambda_{i}.$
From Assumption 2(b) we have
$\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}$ $\overset{p}{\to}$ $S\text{ as }N,T\to\infty$
Further, $\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}\epsilon_{it}=O_{p}\left((NT)^{-1/2}\right)$ as $N,T\to\infty$. Finally, from Assumption 2(c)
$\sum_{i=1}^{N}\sum_{t=1}^{T}x_{it}x_{it}^{\prime}\lambda_{i}=\frac{1}{\gamma}\sum_{i\in G_{2}}^{N}\sum_{t=1}^{T}{\lambda}x_{it}x_{it}^{\prime}=O_{p}\left(\frac{MT}{\gamma}\right),\mbox{as }N,T\to\infty.$
Proof Theorem 1, BRS Test.
We start by considering the $LM$ test of BRS under $H_{1,n}$. After substituting $\hat{\epsilon}_{it}=u_{it}-x_{it}^{\prime}(\hat{\beta}_{LS}-\beta)$, with $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$, into the score vector, we obtain the following decomposition,
$\hat{s}_{k}=\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\epsilon_{it}\epsilon_{is}x_{it,k}x_{is,k}+\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda_{i}^{\prime}{x}_{it}x_{it,k}x_{is,k}{x}_{is}^{\prime}\lambda_{i}+R_{NT,k}+C_{NT,k},$ (A1)
where the first two terms drive the asymptotic normality and the power of the test, respectively, while $R_{NT,k}$ and $C_{NT,k}$ are remainder terms given by,
$R_{NT,k}$ $=$ $\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}(\hat{\beta}_{LS}-\beta)^{\prime}x_{it}x_{it,k}x_{is,k}x_{is}^{\prime}(\hat{\beta}_{LS}-\beta)-2\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}(\hat{\beta}_{LS}-\beta)^{\prime}x_{it}u_{is}x_{it,k}x_{is,k},$
$C_{NT,k}$ $=$ $\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda_{i}^{\prime}{x}_{it}\epsilon_{is}x_{it,k}x_{is,k}+\sum_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\epsilon_{it}\lambda_{i}^{\prime}{x}_{is}x_{it,k}x_{is,k},$
for k=1,,Kk=1,...,K. After substituting $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$ into $R_{NT}$ and using Lemma 1, Lemma A2 in BRS, together with the fact that $\lambda_{i}$ is non zero for only $M$ elements, we obtain,
$\frac{1}{\sqrt{N}T}R_{NT,k}=O_{p}\left(\frac{M^{2}T}{\gamma^{2}{N}^{3/2}}\right).$
Furthermore, if $M/N\to{m_{0}}>0$ then the rate above reduces to $\frac{1}{\sqrt{N}T}R^{(k)}_{NT}=O_{p}\left(\frac{MT}{\gamma^{2}{N}^{1/2}}\right)$ and as $MT/\sqrt{N}\gamma^{2}\to c>0$ we have
$\frac{1}{\sqrt{N}T}R_{NT,k}$ $=c^{2}m_{0}\lambda^{\prime}\left(S^{(G_{2})}S^{-1}\Omega_{k}S^{-1}S^{(G_{2})}-2S^{(G_{2})}S^{-1}\Omega_{k}^{(G_{2})}\right)\lambda+o_{p}(1).$
By the same arguments,
$\frac{1}{\sqrt{N}T}C_{NT,k}=O_{p}\left(\frac{\sqrt{MT}}{\sqrt{N}\gamma}\right),$
and if $MT/(\sqrt{N}\gamma^{2})\to c\geq 0$ then $\frac{\sqrt{MT}}{\sqrt{N}\gamma}\to 0$ resulting in $\frac{1}{\sqrt{N}T}C_{NT,k}=o_{p}(1)$. Further,
$\frac{1}{\sqrt{N}T}\sum\limits_{i=1}^{N}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda_{i}^{\prime}{x}_{it}x_{it,k}x_{is,k}{x}_{is}^{\prime}\lambda_{i}$ $=\frac{1}{\gamma^{2}\sqrt{N}T}\sum_{i\in G_{2}}\sum\limits_{t=2}^{T}\sum\limits_{s=1}^{t-1}\lambda^{\prime}{x}_{it}x_{it,k}x_{is,k}{x}_{is}^{\prime}\lambda$
$\overset{p}{\rightarrow}\lambda^{\prime}\Omega_{k}^{(G_{2})}\lambda\times\lim\limits_{N,T\to\infty}\frac{MT}{\sqrt{N}\gamma^{2}}.$
Let $s_{k}:=\sum_{i=1}^{N}\sum_{t=2}^{T}\sum_{s=1}^{t-1}\epsilon_{it}\epsilon_{is}x_{it,k}x_{is,k}$ and $s=[s_{1},...,s_{K}]^{\prime}$ then by Theorems 5.1 and Corollary 5.2 in BRS,
$\frac{1}{\sqrt{N}T}s\overset{d}{\rightarrow}N(0,\Psi)\text{ as }N,T\to\infty.$
Finally, following Theorem 6.2 in BRS, we can write each $(k,l)$ element of $\hat{V}$ as,
$\frac{1}{NT^{2}}\hat{V}_{k,l}$ $=\frac{1}{NT^{2}}\sum\limits_{i=1}^{N}\sum\limits_{t=2}^{T}\epsilon_{it}^{2}x_{it,k}x_{it,l}\left(\sum\limits_{s=1}^{t-1}\epsilon_{is}x_{is,k}\right)\left(\sum\limits_{p=1}^{t-1}\epsilon_{ip}x_{ip,k}\right)+O_{p}\left(\frac{MT}{\gamma^{2}N}\right)$
$=\Psi_{k,l}+O_{p}\left(\frac{MT}{\gamma^{2}N}\right).$
Putting all these results together into the expression for the LM test (4) yields the first statement of the theorem.
Proof Theorem 1, PY Test.
We now use the same procedure to prove the statement for the PY test. Under $H_{1,n}$, we have that,
$\left(\hat{\beta}_{i}-\hat{\beta}_{WLS}(\tilde{\sigma}_{i})\right)=\left[\lambda_{i}+\left(\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\frac{x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}\right)-\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}u_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right]$
where $u_{i}=(u_{i1},...,u_{iT})^{\prime}=x_{i}\lambda_{i}+\epsilon_{i}$. Thus, rewrite $S_{PY}$ as,
$S_{PY}=\sum_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}+\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}+R_{NT}+C_{NT},$
where,
$R_{NT}$ $=$ $-\left(\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}\right)-\left(\sum\limits_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}\right),$
$C_{NT}$ $=$ $2\sum_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}-2\left(\sum\limits_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)^{-1}\left(\sum\limits_{i=1}^{N}\frac{x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}\right).$
Using that $(NT)^{-1}\sum_{i=1}^{N}(x_{i}^{\prime}x_{i}/\tilde{\sigma}^{2}_{i})=O_{p}(1+T^{-1/2})$ and $(NT)^{-1/2}\sum_{i=1}^{N}(x_{i}^{\prime}\epsilon_{i}/\tilde{\sigma}^{2}_{i})=O_{p}(1+T^{-1/2})$ (see equations A.10, A.11, and Lemma 3 in PY), it follows that,
$\frac{1}{\sqrt{N}}R_{NT}$ $=O_{p}\left(\frac{M^{2}T}{\gamma^{2}{N}^{3/2}}\right).$
Furthermore, if $M/N\to{m_{0}}>0$ then the rate above reduces to $\frac{MT}{\gamma^{2}{N}^{1/2}}$, as $M/N<1$, and if $MT/\sqrt{N}\gamma^{2}\to c>0$ we have
$\frac{1}{\sqrt{N}}R_{NT}=-(c)^{2}m_{0}\lambda^{\prime}Q^{(G_{2})}Q^{-1}Q^{(G_{2})}\lambda.$
By the same arguments
$\frac{1}{\sqrt{N}}C_{NT}$ $=O_{p}\left(\frac{\sqrt{MT}}{\gamma\sqrt{N}}+\frac{M\sqrt{T}}{\gamma N}\right),$
and,
$\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}=\frac{MT}{\sqrt{N}\gamma^{2}}\lambda^{\prime}Q^{(G_{2})}\lambda+O_{p}\left(\frac{M\sqrt{T}}{\sqrt{N}\gamma^{2}}\right).$
Finally, by the same arguments of Theorem 1 and 2 in PY (see appendix A.2, A.3 of PY),
$\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\frac{\epsilon_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\epsilon_{i}}{\tilde{\sigma}^{2}_{i}}=\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\tilde{z}_{i}+O_{p}\left(\frac{1}{\sqrt{T}}\right)+O_{p}\left(\frac{1}{\sqrt{N}}\right),$
where $\tilde{z}_{i}=(T-1)\epsilon_{i}^{\prime}x_{i}(x_{i}^{\prime}x_{i})^{-1}x_{i}^{\prime}\epsilon_{i}/(\epsilon_{i}^{\prime}\epsilon_{i})$ with $\mathbb{E}\left[\tilde{z}_{i}\right]=K+O(T^{-2})$ and $\operatorname{Var}{\tilde{z}_{i}}=2K+O(T^{-1})$ uniformly over $i$. Then putting all these rates together into expression (2) for PY test yields
$\Delta$ $=\sqrt{N}\left(\frac{N^{-1}S_{PY}-K}{\sqrt{2K}}\right)=\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\left(\frac{\tilde{z}_{i}^{*}-K}{\sqrt{2K}}\right)+O\left(\frac{\sqrt{N}}{T^{2}}\right)$
$+\frac{MT}{\sqrt{N}\gamma^{2}}\lambda^{\prime}\left(\frac{1}{MT}\sum_{i\in G_{2}}\frac{x_{i}^{\prime}x_{i}}{\sigma_{i}^{2}}\right)\lambda\frac{1}{\sqrt{2K}}+\frac{MT}{\sqrt{N}\gamma^{2}}O_{p}\left(\frac{M}{N}\right)\frac{1}{\sqrt{2K}}+o_{p}(1),$
where $N^{-1/2}\sum_{i=1}^{N}(\tilde{z}_{i}-K)/\sqrt{2K}\xrightarrow{d}N(0,1)$ (appendix A.4 of PY).
The statement in Theorem 1 follows immediately by considering different scenarios for the rate $\frac{MT}{\sqrt{N}\gamma^{2}}$. ∎
Proof Theorem 1, SC Test.
After substituting $\hat{\epsilon}_{i}=\epsilon_{i}+x_{i}\lambda_{i}-x_{i}(\hat{\beta}_{LS}-\beta)$ into $LM_{SC}-\sqrt{N}\hat{B}_{NT}$ we obtain,
$LM_{SC}-\sqrt{N}\hat{B}_{NT}=\sum_{i=1}^{N}\epsilon_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})\epsilon_{i}+\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}\lambda_{i}+R_{NT}+C_{NT}$
where $\operatorname*{diag}\{A\}$ indicates the square diagonal matrix with main elements given by the diagonal of $A$, and
$R_{NT}$ $=\sum_{i=1}^{N}(\hat{\beta}_{LS}-\beta)^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}(\hat{\beta}_{LS}-\beta)-2\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}(\hat{\beta}_{LS}-\beta),$
$C_{NT}$ $=2\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})\epsilon_{i}-2\sum_{i=1}^{N}\epsilon_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}(\hat{\beta}_{LS}-\beta).$
Using identical arguments to the ones used for BRS test we have,
$\frac{1}{\sqrt{N}}R_{NT}$ $=O_{p}\left(\frac{M^{2}T}{{N}^{3/2}\gamma^{2}}\right),$
where we explicitly used the fact that $h_{i}$ is a projection matrix, hence $0\leq h_{i,tt}\leq 1$, and
$\sum_{i=1}^{N}\lVert x_{i}^{\prime}\left(h_{i}-\operatorname*{diag}\{h_{i}\}\right)x_{i}\rVert_{2}\leq\sum_{i=1}^{N}\sum_{t=1}^{T}\lVert(x_{it}x_{it}^{\prime})(1-h_{i,tt})\rVert_{2}\leq\sum_{i=1}^{N}\sum_{t=1}^{T}\lVert x_{it}x_{it}\rVert_{2}=O_{p}(NT).$
Here $\lVert A\rVert_{2}=\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{k}|a_{ij}|^{2}}$ denotes the Frobenius norm of a $k\times k$ matrix $A$. When $M/N\to{m_{0}}>0$ and $MT/\sqrt{N}\gamma^{2}\to c>0$ then
$\frac{1}{\sqrt{N}}R_{NT}=(c)^{2}m_{0}\lambda^{\prime}\left(S^{(G_{2})}S^{-1}WS^{-1}S^{(G_{2})}-W^{(G_{2})}S^{-1}S^{(G_{2})}-S^{(G_{2})}S^{-1}W^{(G_{2})}\right)\lambda+o_{p}(1).$
Similarly,
$\frac{1}{\sqrt{N}}C_{NT}$ $=O_{p}\left(\frac{\sqrt{MT}}{\sqrt{N}\gamma^{2}}\right),$
and if $MT/(\sqrt{N}\gamma^{2})\to c\geq 0$ then $\frac{1}{\sqrt{N}}C_{NT}=o_{p}(1)$. Further,
$\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\lambda_{i}^{\prime}x_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})x_{i}\lambda_{i}\xrightarrow{p}\lambda^{\prime}W^{(G_{2})}\lambda\times\lim_{N,T\to\infty}\frac{MT}{\sqrt{N}\gamma^{2}}.$
Finally, applying Theorem 5.1 in BRS, under $H_{1,n}$, we obtain $\hat{V}_{NT}=V_{0}+o_{p}(1)$, and
$\sum_{i=1}^{N}\epsilon_{i}^{\prime}(h_{i}-\operatorname*{diag}\{h_{i}\})\epsilon_{i}\to N(0,V_{0}).$
Putting all of the obtained rates into expression (3) of SC test gives us the non-centrality parter and the statement of the theorem. ∎
Remarks Corollary 1.
The term $C_{NT}$ in all decompositions can be neglected under the conditions of Theorem 1 with $M=\mathbb{M}$. We show the passages to obtain $\delta_{PY}$; the remarks for the other two tests are similar and thus omitted.
Let $u_{it}=x_{it}^{\prime}\lambda_{i}+\epsilon_{it}$ where $\lambda_{i}=\sum_{p=2}^{P}\lambda_{p}\mathbb{I}\{g(i)=p\}$, with $\mathbb{I}\{A\}$ being the indicator of event $A$. Then,
$\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\lambda_{i}^{\prime}x_{i}^{\prime}x_{i}\lambda_{i}}{\tilde{\sigma}^{2}_{i}}$ $=\frac{1}{\sqrt{N}\gamma^{2}}\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\lambda_{p}=\sum_{p=2}^{P}O_{p}\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)=O_{p}\left(\frac{\mathbb{M}T}{\sqrt{N}\gamma^{2}}\right),$
with
$\mathop{\rm plim}_{N,T\to\infty}\frac{1}{\sqrt{N}\gamma^{2}}\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\lambda_{p}=\lim_{N,T\to\infty}\sum_{p=2}^{P}\lambda_{p}^{\prime}Q^{(G_{p})}\lambda_{p}\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1).$
Using the same arguments in $R_{NT}$, we obtain,
$\frac{1}{\sqrt{N}}R_{NT}$ $=-\frac{1}{\sqrt{N}\gamma^{2}}\left(\sum_{p=2}^{P}\lambda_{p}^{\prime}\left(\sum_{i\in G_{p}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)\frac{Q^{-1}}{NT}\left(\sum_{q=2}^{P}\lambda_{p}\left(\sum_{i\in G_{q}}\frac{x_{i}^{\prime}x_{i}}{\tilde{\sigma}^{2}_{i}}\right)\right)+o_{p}(1)$
$=\sum_{p=2}^{P}\sum_{q=2}^{P}\lambda_{q}^{\prime}Q^{(G_{p})}Q^{-1}Q^{(G_{q})}\lambda_{q}\left(\frac{M_{p}}{N}\right)\left(\frac{M_{p}T}{\sqrt{N}\gamma^{2}}\right)+o_{p}(1)$
Appendix B Formal Statement Fixed $T$
Assumption 3 (Regressors)
  • (a)
    $\mathbb{E}|x_{it,k}|^{4+\delta}<C<\infty$, for some $\delta>0$, and all i=1,,Ni=1,...,N, t=1,,Tt=1,...,T, and k=1,,Kk=1,...,K.
  • (b)
    For any subset $G\subset\{1,...N\}$ such that $G\to\infty$, the probability limits $\Sigma=\mathop{\rm plim}_{N\to\infty}N^{-1}\sum_{i=1}^{N}x_{i}^{\prime}x_{i}$ and $\Sigma^{(G)}=\mathop{\rm plim}_{N\to\infty}(\#G)^{-1}\sum_{i\in G}x_{i}^{\prime}x_{i}$ are positive definite, non-stochastic and finite matrices.
Using the proof of Theorem 1, we obtain the following result:
Corollary 2
Under Assumptions 1, 2, and $H_{1,n}$, as $M,N\to\infty$, the following holds:
$LM\overset{d}{\rightarrow}\begin{cases}\chi_{K}^{2},&\mbox{if }\frac{\sqrt{M}}{\gamma N^{1/4}}\to 0,\\ \chi_{K}^{2}\left(\delta_{LM}^{\prime}\mathcal{V}^{-1}\delta_{LM}\right),&\mbox{if }\frac{\sqrt{M}}{\gamma N^{1/4}}\to c>0\\ \infty,&\mbox{if }\frac{\sqrt{M}}{\gamma N^{1/4}}\to\infty.\end{cases},$
where $\delta_{LM}$ is a $K\times 1$ vector with $k$-th element
$\delta_{LM,k}=c^{2}\left(\lambda_{2}^{\prime}U_{k}^{(G_{2})}\lambda_{2}+m_{0}\lambda_{2}^{\prime}\left(\Sigma^{(G_{2})}\Sigma^{-1}U_{k}\Sigma^{-1}\Sigma^{(G_{2})}-2\Sigma^{(G_{2})}\Sigma^{-1}U_{k}^{(G_{2})}\right)\lambda_{2}\right),$
with
$U_{k}^{(G)}=\underset{\#G\rightarrow\infty}{\text{plim}}\;(\#G)^{-1}\sum_{i\in G}\sum_{s<t}x_{it,k}x_{it}x_{is,k}x_{is}^{\prime},$
and $U_{k}=U_{k}^{(G_{1}\cup G_{2})}$. Here, $\mathcal{V}$ is a $K\times K$ matrix with $\left(k,l\right)$-th element given by $\mathcal{V}_{k,l}=\underset{N\rightarrow\infty}{\text{plim}}\;N^{-1}\sum_{i=1}^{N}\sum_{s<t}\sigma_{i}^{4}x_{it,k}x_{it,l}x_{is,k}x_{is,l}$.